### v1

# Are foundation models for computer vision good conformal predictors?

Abstract
Recent advances in self-supervision and constrastive learning have brought the performance of foundation models to unprecedented levels in a variety of tasks. Fueled by this progress, these models are becoming the prevailing approach for a wide array of real-world vision problems, including risk-sensitive and high-stakes applications. However, ensuring safe deployment in these scenarios requires a more comprehensive understanding of their uncertainty modeling capabilities, which has been barely explored. In this work, we delve into the behavior of vision and vision-language foundation models under Conformal Prediction (CP), a statistical framework that provides theoretical guarantees of marginal coverage of the true class. Across extensive experiments including popular vision classification benchmarks, well-known foundation vision models, and three CP methods, our findings reveal that foundation models are well-suited for conformalization procedures, particularly those integrating Vision Transformers. Furthermore, we show that calibrating the confidence predictions of these models leads to efficiency degradation of the conformal set on adaptive CP methods. In contrast, few-shot adaptation to downstream tasks generally enhances conformal scores, where we identify Adapters as a better conformable alternative compared to Prompt Learning strategies. Our empirical study identifies APS as particularly promising in the context of vision foundation models, as it does not violate the marginal coverage property across multiple challenging, yet realistic scenarios.
1 Introduction
Large-scale pre-trained vision foundation models, such as DINOv2 [33], as well as those integrating text, such as CLIP [38], are driving a new learning paradigm in machine learning, achieving unprecedented results on a broad spectrum of tasks. Nevertheless, despite their desirable zero-shot and generalization capabilities, recent evidence has pointed out to the existence of bias and factual errors in these models [49], which transcend the field of computer vision [44, 4]. For example, the original CLIP paper [38] demonstrated gender and race biases in certain zero-shot tasks, whereas [49] identified that CLIP-based models are not always better calibrated than other arguably simpler ImageNet models. Furthermore, [30] recently showed that adapted models magnify the miscalibration issue compared to the zero-shot setting, yielding overconfident predictions. These problems underscore widespread societal concerns surrounding the reliable deployment and use of foundation models in sensitive contexts, such as decision-making processes in critical scenarios, e.g., healthcare, autonomous driving or security applications.
A popular solution to quantify the uncertainty present in the predictions of deep models is calibration. In this setting, the proposed strategies aim at reducing the discrepancies between model predictions and the actual correctness probability. Temperature Scaling (TS) [13], a simple variant of Platts Scaling [36], provides a simple post-processing approach to adjust the softmax probability scores of the trained models. Other line of methods have proposed training objectives to enforce the model to produce less confident scores, either in the predictions space [37, 35], logits [27, 26], or modifying the ground truth labels [28, 29]. Nevertheless, confidence calibration does not offer theoretical guarantees for coverage of the true class.
In contrast, Conformal Prediction (CP) [51], an alternative strategy to quantify the uncertainty [41, 40, 1], is a statistical framework which offers several advantages over calibration methods. First, unlike most methods in calibration, CP works directly on the model predictions, offering an appealing solution for black-box models. Second, instead of simply modeling the correctness of the predicted probability, CP produces a set of predictions, including the most likely classes, which can be of much interest in certain problems. Last, CP methods have theoretical guarantees for the marginal coverage of the true class within the predicted set, under several assumptions.
Due to these properties, CP is gaining attention within the community to conformalize the predictions of deep models [2, 1, 9, 42]. Nevertheless, albeit the efforts to study CP in large language foundation models [12], its impact on vision foundation models has been unexplored, besides the significant implications it may have on a variety of vision problems. The aim of this study is to shed light and provide important insights into this direction. To achieve this, we conducted an extensive empirical analysis of the performance of three common CP methods on ten popular vision foundation models across three vision datasets. Our extensive experiments further explore common situations encountered in practice, assessing the impact on CP: under distributional drifts, after confidence calibration and in few-shot adaptation to novel downstream tasks. Our key observations are:
  1. (i)
    Vision, and vision-language foundation models seem to better conformalize compared to their more traditionally (fully-supervised) trained counterparts.
  2. (ii)
    Across different foundation models, those including visual transformers, such as DINO and CLIP, lead to better conformal results compared to models integrating convolutional neural networks.
  3. (iii)
    Across all the experiments, Adaptive Prediction Sets (APS) is the best Conformal Prediction approach in terms of empirical coverage, while Regularized Adaptive Prediction Sets (RAPS) presents the best alternative from a conformal set size standpoint.
  4. (iv)
    Under distributional shifts, APS exhibits the highest robustness among CP methods in terms of coverage guarantes, albeit decreasing its set efficiency.
  5. (v)
    Confidence calibration decreases the efficiency of conformal sets, but typically improves coverage gap.
  6. (vi)
    Adapters are generally better candidates for CP than Prompt Learning methods, in both in-distribution and out-of-distribution scenarios.
2 Related Work
Foundation models for computer vision.  The landscape of foundation models has rapidly evolved in the last years. Traditionally, pretrained convolutional networks based on ResNet architectures [15] were the main models used by the community. However, recently driven by the unprecedented advances in language models, such as GPT [6] or LLaMA [48], as well as the vast availability of image data online, there are groundbreaking advances in unimodal [7, 33, 22] and multimodal [38] foundation models for vision tasks based mainly on vision transformer architectures. These large pre-trained models aim to generalize across a broad range of visual tasks by pre-training on massive, diverse image datasets, exhibiting strong zero-shot and generalization capabilities to new tasks. For example, vision foundation models, such as DINO [7, 33], rely on self-supervised learning strategies on large datasets, which results in an excellent semantic understanding of the visual content. On the other hand, CLIP [38] bridges the gap between language and vision modalities through contrastive learning, effectively allowing the model to understand images in the context of natural language prompts and enabling zero-shot capabilities.
Uncertainty estimation in deep models.  Quantifying the uncertainty of the predictions of deep networks has recently garnered considerable interest. From a calibration standpoint, popular strategies include post-hoc approaches [13, 10, 20], which map the logits or softmax predictions to smoother distributions, and explicit learning objectives [26, 27, 29, 28, 16, 31], which are integrated into the loss function. Nevertheless, a main limitation of calibration methods is that they lack theoretical guarantees of model performance. In contrast, CP has recently emerged as a promising alternative, which provides marginal coverage guarantees over unseen test samples [52, 43]. Specifically, CP resorts to a non-conformity score function (i.e., a measure of how “different” a particular data point is compared to a CP calibration dataset) to produce a finite prediction set, which is guaranteed to contain the true label with a specified confidence level. A central objective of the CP literature has been to improve either the set efficiency (i.e., smaller set sizes) or the class conditional coverage. For this purpose, several non-conformity scores have been presented in the literature [47, 40, 1, 41, 11, 9, 46], with [40, 1, 41] being popular methods widely studied. A straighforward solution consists on directly using the raw class softmax predictions to generate the prediction sets [41]. Adaptive Prediction Sets (APS) [40] provides an adaptive version, by computing the non-conformity scores by accumulating sorted softmax probabilities in descending order. To further improve the efficieny, RAPS [1] introduces an explicit regularization term, which penalizes non-conformity scores for unlikely classes.
However, a main limitation of existing evaluations is the focus on more traditional models, usually trained on a data collection that falls within the distribution of calibration and test data points. Thus, given the rising popularity of foundation models in strategic domains, quantifying the uncertainty of their predictions is paramount. However, whereas uncertainty quantification from a calibration perspective has been scarcely studied [30, 55, 50], its exploration under CP is, to the best of our knowledge, overlooked.
3 Background
3.1 Conformal Prediction Framework
Let $\mathcal{X}$ and $\mathcal{Y}$ denote the input and output space, respectively. We assume access to a calibration set $\mathcal{D}_{cal}=\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{N}$ of $N$ independent and identically distributed (i.i.d.) samples, where each $\mathbf{x}_{i}=(p_{ik})_{1\leq k\leq K}$ represents the black-box probabilities and $y_{i}\in\mathcal{Y}=\{1,2,...,K\}$ is the associated label. The goal of CP is to construct a prediction set $\hat{C}(\mathbf{x}_{n+1})\subseteq\mathcal{Y}$ for a new test input $\mathbf{x}_{n+1}$ such that it contains the true label $y_{n+1}$ with a user pre-specified coverage probability $1-\alpha$, where $\alpha\in(0,1)$ denotes the error level.
The core idea of CP is to assess the degree to which a new sample conforms to the underlying distribution of the calibration data by computing non-conformity scores. Let $S(\mathbf{x},y)$ be a non-conformity measure (or scoring function) that assigns a score $s_{i}$ to each $(\mathbf{x}_{i},y)$ pair, quantifying how unusual the pair is relative to the rest of the data. Given the non-conformity scores for all calibration examples and a new input $\mathbf{x}_{n+1}$ (unseen in the calibration set), the conformal prediction set $C(\mathbf{x}_{n+1})$ is defined as:
$C(\mathbf{x}_{n+1})=\left\{y\in\mathcal{Y}:S(\mathbf{x}_{n+1},y)\leq q_{\alpha}\right\},$ (1)
where $q_{\alpha}$ is the $1-\alpha$ quantile of the non-conformity scores on the calibration set, obtained with the observed labels:
$q_{\alpha}=\textsc{Quantile}\Big{(}\{S(\mathbf{x}_{i},y_{i})\}_{i=1}^{N},\frac{\lceil(n+1)(1-\alpha)\rceil}{n}\Big{)}$ (2)
Coverage Guarantees.  A key property of conformal prediction is its finite-sample coverage guarantee. This property ensures that the prediction sets achieve the desired coverage probability marginally over $\mathcal{X}$ and $\mathcal{Y}$, irrespective of the underlying data distribution, as long as the calibration and test data are exchangeable [52]. Formally, for any $1-\alpha$, conformal predictors satisfy:
$\mathbb{P}(y_{n+1}\in C(\mathbf{x}_{n+1}))\geq 1-\alpha.$ (3)
This property is crucial for applications requiring reliable uncertainty quantification, particularly where distributional assumptions (e.g., Gaussianity) may not hold.
Tightness of Prediction Sets.  Conformal prediction guarantees valid marginal coverage. Nevertheless, the efficiency of the prediction sets, i.e., their size, depends on the choice of the non-conformity measure. Choosing an appropriate $S(\mathbf{x},y)$ is key to balancing the trade-off between coverage and tightness of the prediction sets. In practice, we aim to minimize the size of the prediction sets while maintaining the desired coverage probability.
3.2 Non-conformity scores
Least Ambiguous Classifier.  LAC [41] aims to construct the smallest possible set under the assumption that the output is correct. Intuitively, it can be interpreted as a thresholding of the output probabilities for each category. Thus, the non-conformity score can be constructed as:
$\mathcal{S}_{\text{\tiny{LAC}}}(\mathbf{x},y)=1-x_{k=y}.$ (4)
LAC also provides notable efficiency in scenarios using an imperfect classifier. However, it lacks adaptability, e.g., in under-represented categories or uncertain predictions.
Adaptive Prediction Sets.  APS [40] provides a non-conformity score that leverages the accumulated confidence in the ordered probability predictions. Thus, APS is known to be an adaptive score, whose main objective is enhancing the coverage of uncertain predictions by sacrificing efficiency. Formally, APS is expressed as:
$\mathcal{S}_{\text{\tiny{APS}}}(\mathbf{x},y)=\rho(\mathbf{x},y)+x_{k=y}\cdot u,$ (5)
where $\rho(\mathbf{x},y)$ is the accumulated confidence of the categories more likely than the evaluated label $y$, i.e., $\rho(\mathbf{x},y)=\sum_{k^{\prime}\in\mathcal{Y^{\prime}}(\mathbf{x},y)}x_{k=k^{\prime}}$, with $\mathcal{Y^{\prime}}(\mathbf{x},y)=\{k|x_{k}>x_{k=y}\}$. Adaptive methods usually include $u\in\{0,1\}$, as a random variable to break ties to archive exact marginal coverage.
Regularized Adaptive Prediction Sets.  RAPS [1] builds upon APS by adding a regularization term to enforce smaller predicted sets. Thus, APS score is modified to penalize the confidence of introducing additional, unlikely categories, after a certain set size is met:
$\mathcal{S}_{\text{\tiny{RAPS}}}(\mathbf{x},y)=\rho(\mathbf{x},y)+x_{k=y}\cdot u+\lambda(o(\mathbf{x},y)-k_{reg})^{+}$ (6)
where $\lambda,k_{reg}\geq 0$ are hyper-parameters controlling the penalty strength, $o_{x}(y)$ is the rank of the sorted label, $o(\mathbf{x},y)=|\mathcal{Y^{\prime}}(\mathbf{x},y)|+1$, and $(\cdot)^{+}$ denotes the positive part.
4 Experiments
4.1 Experimental Setup
Models: We employ a total of ten foundation models: two DINO [7] (DINO-S and DINO-B), four DINOv2 [33] (DINOv2-S, DINOv2-B, DINOv2-L, and DINOv2-G), three VICReg [3] (based on ResNet-50, ResNet-50x2, and ResNet-200x2), and one vision-language model (CLIP [38] based on ViT-B). Our analysis is conducted on three popular vision datasets: CIFAR-10 [23], CIFAR-100 [23], and Imagenet [8], including its versions integrating domain shifts [17, 18, 53, 39]. Each dataset is split into two sets: one for training the linear probing head, and one for the conformal experiments. The latter is then split into one calibration set to tune the CP method, and one test set for evaluation.
Metrics: We resort to conformal set size (a.k.a efficiency), minimum class-conditional coverage (MCCC), coverage gap (CovGap) and marginal coverage to benchmark the different methods. More details about metrics and implementation details can be found in Appendix A and Appendix B.
Figure 1: Relationship between the linear probing model accuracy and conformal set size (top) and the coverage gap (bottom) across different tasks of increasing complexity. From left to right: CIFAR-10, CIFAR-100 (middle) and ImageNet.
Adaptation to target tasks.  The foundation models used in this study have been pre-trained using different strategies. Nevertheless, they need to be adapted for novel tasks, as their pre-trained versions do no accommodate classification tasks, i.e., there is no classification head. To do this, foundation models are frozen and a Linear Probing head (one linear layer followed by a softmax activation function) is trained on each dataset by optimizing a cross-entropy loss.
4.2 Results
To gain insights into the factors influencing the efficacy of CP approaches in vision foundation models, we design four experiments. First, we start by exploring the impact of CP in standard scenarios, where a large calibration set is available to conformalize the predictions of different models. Then, we challenge the status-quo of CP, and alter the conformal sets to accommodate real-world scenarios, by including domain shifts. Furthermore, the notion of confidence calibration being very linked to CP, we explore the impact of calibrating a model on the performance of CP. Last, we explore the impact of CP when adapting a very popular VLM, i.e., CLIP, to novel tasks.
4.2.1 Performance in the General Setting
First, we study the performance of ten vision foundation models paired with CP methods under the standard setting, and on the three datasets, which presents ideal conditions: a sufficiently large calibration set, and absence of distributional drifts between calibration and test sets. Our goal is to determine whether we can prescribe a winner solution in this scenario, and which factors can help its identification.
Fig. 1 depicts the relationship between the linear probing performance of each model and the conformal set size (top), and coverage gap (bottom). The initial observation highlights a clear trend: models with higher performance tend to produce smaller prediction sets, regardless of the conformal method used. Nevertheless, while set efficiency (i.e. size) is typically considered as a sufficient condition in most prior literature in conformal prediction [1], our analysis revealed that the relation with the accuracy does not consistently hold when examining other metrics. In particular, Fig. 1 (bottom) exposes that different CP methods yield mixed results for the coverage gap, which do not correlate directly with accuracy across all CP approaches. While APS appears to be almost unaffected by the model performance, RAPS clearly benefits from a strong performance, and LAC is negatively affected by more accurate models as the dataset becomes more complex.
In terms of comparison between the CP methods, if we consider the set size, APS is clearly outperformed by the other approaches, particularly on ImageNet, whereas LAC provides the smallest prediction sizes, closely followed by RAPS. Indeed, RAPS is specifically designed to reduce the conformal set size of APS. However, the coverage gap results indicate that this comes at the cost of increasing the range of the class-conditional coverage. Below, we analyze the underlying causes that may explain this behaviour.
RAPS class-conditional coverage, and therefore coverage gap, are more sensitive to the accuracy of the model. Let us assume we have two models, $M_{1}$ and $M_{2}$, in a multi-class classification problem, whose accuracies are $Acc_{1}$ and $Acc_{2}$, respectively. For each class $y\in\mathcal{Y}$, we refer to $C_{M_{i}}(y)=\mathbb{P}(Y=y|Y\in\mathcal{S}_{M_{i}}(X))$ as the class-conditional coverage for $y$ under model $M_{i}$, which measures the probability that predictions include the true label when the true label is $y$. Furthermore, let $\delta_{M_{1}}=\min_{y\in\mathcal{Y}}C_{M_{1}}(y)$ and $\delta_{M_{2}}=\min_{y\in\mathcal{Y}}C_{M_{2}}(y)$ denote the minimum class-conditional coverage achieved by each model. Under this scenario, we argue that due to the penalty in RAPS, if $Acc_{1}<Acc_{2}$, then $\delta_{M_{1}}<\delta_{M_{2}}$ when using RAPS as conformal prediction method. In particular, $M_{1}$ (with lower accuracy) needs to expand its prediction sets for certain classes in order to meet the marginal coverage target 1-$\alpha$. Nevertheless, the penalty term encourages small prediction sets, limiting an excessive number of classes. Thus, for some difficult classes, model $M_{1}$ may still potentially fail to meet the target coverage, as the enforced penalty discourages overly large sets. This ultimately results in lower coverage rates for those specific classes compared to $M_{2}$. In contrast, since APS does not include any regularization term that encourages small set sizes, it will compensate for lower performing models by increasing its set sizes, ultimately attaining higher class-conditional coverages.
Figure 2: Comparison (APS vs RAPS) of the class-conditional coverage and set size for the class for which RAPS has the worst class-conditional coverage. Experiments performed on CIFAR-100. Models sorted (in ascending order) by their LP performance ($\min=$0.65 and $\max=$0.92), indicated by the size of the circles.
To confirm this hypothesis, we perform the following experiment across all foundation models. First, we identify the class with the lowest minimum class-conditional coverage obtained by RAPS for each model $m$, which we refer to as $k^{R}_{m}$, and find its corresponding set size (both values are represented by pink circles in Fig. 2). Then, we identify the class-conditional coverage and set size of the class $k^{R}_{m}$ provided by APS across all models, whose values are shown as blue circles in both plots of Fig. 2. Last, we also include the minimum class-conditional coverage obtained by APS across each method, depicted with a cross. Note that the minimum class-conditional coverage from APS does not necessarily correspond to the coverage of class $k^{R}_{m}$. Upon close examination of these results, we observe that, indeed, for models presenting lower accuracy (i.e., those with smaller circles), the gap between the class-conditional coverage for the worst class in RAPS and the same class in APS is consistently larger than for more accurate models (larger circles).
This empirical evidence shows that, because RAPS is constrained from expanding its predictions set size, the minimum class-conditional coverage is substantially reduced. This effect is more pronounced in less accurate models, where the true class may rank far away from the maximum allowable set size in the softmax predictions. In contrast, APS tolerates better worse models by increasing the set size, which ultimately reduces the efficiency of the sets but yields better class-conditional coverages. Last, LAC presents structural differences with RAPS and APS, as it lacks an adaptive mechanism and instead relies on a uniform fixed threshold. Thus, LAC may yield inconsistent coverage rates across classes, resulting in high variability in the class-conditional coverage and thus in the coverage gap.
Set size ($\downarrow$) MCCC ($\uparrow$)
Acc ($\uparrow$) Thr APS RAPS Thr APS RAPS
$\text{ViT}_{\text{CLIP}}$ 72.01 3.03 9.50 3.73 0.434 0.556 0.418
$\text{ViT}_{\text{ImageNet}}$ 76.08 2.36 38.75 4.46 0.416 0.495 0.405
Table 1: Results on ImageNet obtained by CLIP (based on ViT-B) and a ViT-B trained in a supervised manner on ImageNet, across accuracy and min class-conditional coverage (MCCC).
Are foundation models better conformable networks? Following this analysis, we are also interested in determining whether a network pre-trained following a more traditional approach (i.e., standard supervised fine-tuning) offers similar conformal capabilities to self-supervised ones. In particular, we select a ViT-B pre-trained on ImageNet, which is the same architecture as the visual encoder of the CLIP model, so the potential differences may come from the different learning strategies, as well as the amount of employed data. The results from this analysis (Table 1) reveal that, despite obtaining lower classification accuracy when using LP on the $\text{ViT}_{\text{CLIP}}$ model, CP methods typically yield to better performance than in $\text{ViT}_{\text{ImageNet}}$. These differences are particularly important under the APS approach, where the set size is significantly degraded on $\text{ViT}_{\text{ImageNet}}$. Moreover, the class-conditional coverage is also substantially affected, particularly for APS, with nearly 6% decrease.
Figure 3: $\text{ViT}_{\text{ImageNet}}$ vs $\text{ViT}_{\text{CLIP}}$. Analyzing the difference in set size between a $\text{ViT}_{\text{CLIP}}$ and $\text{ViT}_{\text{ImageNet}}$. Equal set sizes not shown.
Figure 4: Evaluation under domain-shift. Set size($\downarrow$), coverage($\uparrow$), and MCCC($\uparrow$) across three CP methods and three foundation models. ImageNet versions are sorted based on their OOD performance in [45].
To further delve into these differences, we compute, for each test sample, the difference between the conformal set size when applied to $\text{ViT}_{\text{ImageNet}}$ and $\text{ViT}_{\text{CLIP}}$ models, whose distribution is depicted in Fig. 3. These values confirm that set size differences are not derived from a small set of isolated outliers, but a considerably large group of samples which see their conformal set increase when using the ViT trained in a supervised manner. These results suggest that the strategies used to train foundation models tend to yield better CP metrics, resulting in conformalized models that can be deployed more safely on critical scenarios.
4.2.2 Impact under Distribution Shifts
The theoretical guarantees of the coverage for conformal prediction hold under the hypothesis that the calibration set and the test set are drawn from the same distribution, i.e., data exchangeability assumption. Thus, in this section, we analyze the impact of having calibration sets that present distributional shifts with regard to the testing set.
We resort to ImageNet and its different versions: ImageNet-R [17], ImageNet-A [18], ImageNet-Sketch [53] and ImageNet-V2 [39]. To introduce the distributional drift between the calibration and testing data, we adapt the pre-trained model to one of the ImageNet versions. Then, ImageNet is used as the calibration set to conform the model, which is later tested on the ImageNet version used for adaptation. This is repeated for each ImageNet variant.
APS [40] exhibits strong robustness against large distributional shifts, at the cost of substantially degrading efficiency. One would expect that adaptive CP methods, such as APS and RAPS, somehow mitigate domain shifts due to their adaptive nature. Nevertheless, Fig. 4 reveals several interesting observations, which contradict this intuition. First, we can observe that, when resorting to APS as CP method the coverage gap is consistently satisfied (or nearly satisfied) across all domain shifts and models (Fig. 4, middle). In contrast, RAPS generally shows very similar performances compared to the straighforward LAC, obtaining lower marginal coverage under several models and domains, and substantially lower than APS. To understand this phenomenon, we now study how the set sizes evolve across domains for the different methods (Fig. 4, left). We can easily observe that (i) APS yield the largest conformal sets across ImageNet domains, regardless of the model, and (ii) APS experiences the largest set increases when the complexity of the domain grows. Thus, as exposed in the previous Section, APS satisfied the marginal coverage by substantially including more predicted classes, therefore increasing conformal set sizes.
Figure 5: Domain shift analysis. Distribution of class-conditional coverages for CLIP on ImageNet-A: APS (left) and RAPS (right).
ECE ($\downarrow$) AvgSize ($\downarrow$) MCCC ($\uparrow$)
(×$\times 10^{-2}$) LAC APS RAPS LAC APS RAPS
DINOv2-S 2.71/1.37 1.87/1.87 6.66/8.34 2.12/2.19 0.370/0.351 0.520/0.538 0.384/0.409
DINOv2-B 2.85/1.81 1.36/1.36 7.50/10.46 1.67/1.77 0.414/0.406 0.476/0.482 0.489/0.498
DINOv2-L 2.99/1.81 1.21/1.21 6.77/9.67 1.50/1.59 0.326/0.336 0.502/0.522 0.453/0.487
DINOv2-G 3.66/2.10 1.18/1.18 4.08/5.69 1.44/1.50 0.244/0.243 0.458/0.484 0.350/0.375
VICReg (RN-50x2) 2.34/2.16 3.08/3.09 13.49/16.06 3.89/3.80 0.442/0.430 0.549/0.557 0.445/0.446
VICReg (RN-100x2) 2.21/1.90 2.50/2.49 12.05/14.61 2.98/3.01 0.440/0.430 0.567/0.566 0.462/0.471
CLIP 2.70/1.63 3.03/3.01 9.50/11.11 3.73/3.53 0.434/0.441 0.556/0.564 0.418/0.414
Table 2: Quantitative impact of calibrating vision foundation models. Average size and minimum class conditional coverage are reported between uncalibrated/calibrated models ($T=1$ and $T=1.1$ respectively) for several models, with ImageNet as benchmark dataset. Arrows indicate a decrease () or an increase() in metric after calibration, color indicates better (green) and worse (red) performance.
Figure 6: Impact of the temperature $T$ on ECE ($\downarrow$), set size ($\downarrow$), MCCC ($\uparrow$), and CovGap ($\downarrow$) (CLIP on ImageNet). Green region indicates the area with ECE smaller than for $T=1$. Green dotted line indicates value of optimal calibration ($T=1.1$). More plots in Appendix D.
APS and RAPS are adaptive methods that produce similar minimum class-conditional coverage, as exhibited in ImageNet-A (Fig. 4, right). However, they showcase significant differences in the average coverage gap. In the following, we explore this phenomenon in detail. More concretely, we depict in Fig. 5 the distribution of the conditional class coverage values obtained by APS and RAPS on ImageNet-A. Interestingly, while both approaches see their minimum class-conditional coverage decrease, their distributions are completely different. Indeed, APS distributions exhibit a Gaussian shape, with a decreasing number of categories presenting lower conditional coverage as they separate from 1-$\alpha$. In contrast, the distribution in RAPS exposes a significantly worse scenario, where class-conditional coverages are almost uniformly spread, with a non-negligible amount of classes below the expected coverage, i.e., $1-\alpha$.
Another interesting, and valuable observation is that the low performance of all CP methods is magnified when coupled with VICReg (i.e., larger set sizes and lower coverage), suggesting that this family of foundation models may suffer more under distribution shifts.
4.2.3 Does calibration affect Conformal Prediction?
Confidence calibration is a popular strategy to improve the uncertainty estimates of deep models. These techniques, which can be either added as a post-processing step [36, 13, 24] or integrated as a training regularizer [16, 5, 32], adapt the model softmax predictions to reflect their actual performance accurately. The exponentially growing adoption of vision and vision-language foundation models in critical areas makes integrating confidence calibration a natural progression, as evidenced by recent works [30, 55, 50]. Thus, in this section, we investigate this important issue, as the relationship between calibration and conformal prediction in vision foundation models remains largely unexplored. Specifically, we examine whether calibrating these models with the popular Temperature Scaling (TS) [13] affects the conformal performance of fixed and adaptive CP methods. Particularly, we apply TS to the ImageNet results in the general case (Section 4.2.1). As a proper validation set is not available, we evaluate the CP performance over a set of $T$ values (14 values from $0.85$ to $2$), and found that $T=1.1$ typically yield well-calibrated models [^1] Note that our goal is not to obtain the best-calibrated model but evaluate the impact of calibration on the conformal predictions..
Calibration decreases the efficiency of conformal prediction sets but increases the minimum class-conditional coverage, improving the coverage gap. Confidence calibration typically smooths the distribution of the class softmax scores, which results in less confident predictions. Consequently, the dominant value in the predicted softmax vector is typically lower in calibrated models. Nevertheless, since CP methods provide theoretical guarantees (under the data exchangeability assumption) [52] to satisfy the target $1-\alpha$ of marginal coverage, these changes in the softmax distributions affect the conformalization obtained pre-TS. In particular, in Table 2, we present the results for the average set size and minimum class-conditional coverage before and after scaling the logits with TS ($T=1.1$). Furthermore, we include the ECE values to verify that model calibration has improved. These results show that if the model is calibrated, its prediction set size tends to be larger, particularly for adaptive CP approaches (RAPS, and more specifically, APS). Fig. 7 further delves into these results, where we plot the distribution of differences between the set size of samples before and after calibration (i.e., a point in the distribution is $\mathcal{C}(x_{i})-\mathcal{C}(x_{i}^{TS})$). We can identify that the overall larger size in APS is caused by a consistent efficiency degradation across samples and not a few atypical cases with large conformal sets.
Figure 7: Difference in conformal set size (i.e., efficiency) when applying temperature scaling on APS and DINOv2-B ($T=1.1$).
We observe that class-conditional coverage is typically improved on calibrated models, particularly when they are conformalized by adaptive CP methods. Despite being marginal, these improvements suggest that calibrating vision foundation models decrease efficiency while marginally enhancing class-conditional coverage.
The best coverage gap obtained by APS [40] approximately aligns with the coverage gap achieved at the optimal calibration point, whereas RAPS [1] coverage gap strongly differs. Fig. 6 shows the impact of the temperature $T$ on the different CP metrics on CLIP conformalized predictions. An interesting observation is that at the optimal temperature value $T=1.1$, the CovGap of APS is very close to its optimal point [^2] Results on Appendix D show similar behavior for other models. (0.0567 vs 0.0561). In contrast, while the RAPS coverage gap at optimal calibration is 0.0701, it decreases to 0.0642 as $T$ increases. This suggests that we could anticipate the best coverage gap for APS by assessing the model calibration performance, which will likely be near optimal calibration. On the downside, APS conformal sets efficiency is degraded. As $T$ increases, the softmax distributions get smoother, explaining the degradation in efficiency, which is even more drastic for APS, whose set size monotonically increases with $T$. We concede that, whereas APS minimizes the coverage gap, it does so at the expense of increasing the conformal set size, particularly compared to other methods. However, we believe that increasing the size, up to some extent, while maintaining a better coverage gap is preferable in critical decision systems.
4.2.4 Effect on few-shot adapted models
Adapting zero-shot CLIP models for downstream tasks in a few-shot labeled regime is becoming increasingly popular. These strategies can be mainly categorized into prompt learning [58, 14, 57], which optimize the set of text prompts given to the text encoder, and Adapters [56, 25, 19, 45, 58, 14], where only a limited set of learnable parameters atop embedding representations is updated. Thus, questioning whether adapting these models hinges the performance of CP methods is of paramount importance, as it addresses a foundational aspect of effectively integrating uncertainty quantification with modern architectures. To assess the impact of CLIP adaptation in conformal prediction, we resort to representative methods of the few-shot adaptation families presented above, and follow standard adaptation and evaluation protocols in the literature [45]: adapt CLIP on few-shot samples ($M=16$) from ImageNet, and evaluate on ImageNet and its variants.
Set size ($\downarrow$) CovGap ($\downarrow$)
LAC APS RAPS LAC APS RAPS
ID ZS 2.81 10.07 3.23 0.086 0.069 0.084
ZSLP 2.17 6.60 2.45 0.076 0.060 0.074
CLAP [45] 2.15 6.83 2.45 0.077 0.060 0.074
CoOp [58] 2.44 6.32 2.79 0.080 0.066 0.080
KgCoOp [14] 2.47 7.18 2.84 0.081 0.065 0.081
OOD ZS 7.68 19.22 9.93 0.095 0.085 0.094
ZSLP 8.34 18.51 10.61 0.092 0.083 0.092
CLAP [45] 7.55 17.41 9.89 0.093 0.083 0.091
CoOp [58] 7.86 16.35 10.03 0.093 0.085 0.095
KgCoOp [14] 7.54 16.68 9.79 0.094 0.085 0.095
Table 3: Set size and CovGap for CLIP (ViT-B backbone) on in-distribution (ImageNet) and out-of-distribution (averaged over ImageNet versions) across 5 few-shot adaptation approaches.
Adapters are generally better candidates for Conformal Prediction than Prompt Learning methods. Table 3 reports the results of adapting CLIP with different approaches. First, we observe that, across both ID and OOD scenarios, Adapters (i.e., ZSLP and CLAP) typically yield smaller, or comparable, conformal set sizes than prompt learning (i.e., CoOp and KgCoOp), while always providing a lower CovGap, even compared to the ZS predictions. Furthermore, and more interestingly, conformalizing the predictions of all adaptation methods with APS in the OOD scenario improves the set efficiency compared to the ZS baseline. In contrast, LAC and RAPS typically increase the set size. These observations are related to the findings from the previous section, which suggested that better-calibrated models lead to larger conformal sets. Indeed, recent evidence [30] demonstrated that few-shot CLIP adaptation methods deteriorate the confidence estimates compared to ZS predictions, which, following our observations, should result in smaller conformal sets, as validated in Table 3.
5 Conclusion
In this study, we aim at answering the question: Which CP method and model should I use, and what can I expect, in the era of vision foundation models? Our findings revealed that vision foundation models yield better conformal scores than their traditional pre-trained counterparts, with models integrating visual transformers outperforming those that use convolutional neural networks, particularly under domain shifts. Furthermore, we observed that several common situations encountered in practice (i.e., presence of distributional drifts and models undergoing confidence calibration) are indeed detrimental for some CP approaches. Interestingly, an adaptive CP method, i.e., APS, exhibited stronger robustness to these scenarios, particularly in terms of conditional coverage, but at the expense of degrading set efficiency. Last, under the popular few-shot CLIP adaptation setting, adapters have demonstrated to be a more promising alternative than Prompt Learning.
The final decision ultimately hinges on the requirements of each task. In certain fields, e.g., medical diagnosis, it may be preferable to maximize the conditional coverage across classes, even if it increases the conformal set size, as this minimizes the risk of critical errors. In contrast, in domains where errors are less consequential, maintaining smaller set sizes might be preferable to streamline decision-making and reduce computational demands. Thus, the choice depends on balancing the trade-offs between accuracy, usability, and the specific demands of the application.
Acknowledgments
This work has benefited from state financial aid, managed by the Agence Nationale de Recherche under the investment program integrated into France 2030, project reference ANR-21-RHUS-0003. This work was partially supported by the ANR Hagnodice ANR-21-CE45-0007. This work was partially supported by the ANR Hagnodice ANR-23-CE45-0029. This work was granted access to the HPC resources of IDRIS under the allocation 2023-AD011014802 made by GENCI. We gratefully acknowledge the DATAIA program for supporting JD as a visiting professor at Université Paris-Saclay.
Supplementary Material
A Metrics
Given a pre-specified coverge probability $1-\alpha$, a test set $\mathcal{D}_{\text{test}}$ and a method for predicting conformal sets $\hat{C}(\cdot)$, we resort to the following metrics.
Set $\text{Set size}(\mathcal{D}_{\text{test}})=\frac{1}{|\mathcal{D}_{\text{test}}|}\sum_{x\in\mathcal{D}_{\text{test}}}|\hat{C}(x)|$ (7)
$\text{Cov}(\mathcal{D}_{\text{test}})=\frac{1}{|\mathcal{D}_{\text{test}}|}\sum_{x,y\in\mathcal{D}_{\text{test}}}\mathbbm{1}_{y\in\hat{C}(x)}$ (8)
$\text{CovGap}(\mathcal{D}_{\text{test}})=\frac{1}{|\mathcal{Y}|}\sum_{k\in\mathcal{Y}}\big{|}\text{Cov}(\mathcal{D}_{\text{test},k})-(1-\alpha)\big{|}$ (9)
$\text{MCCC}(D_{\text{test}})=\min_{k\in\mathcal{Y}}\big{(}\text{Cov}(\mathcal{D}_{\text{test},k})\big{)}$ (10)
Where Equation 7 is the average predicted set size across the test set (sometimes refered to as efficiency), Equation 8 is the empirical marginal coverage, Equation 9 represents the average gap between the empirical coverage and the class-conditioned coverage, and Equation 10 is the min class-conditioned coverage (MCCC).
B Models Used and Implementation Details
Foundation models.
We employ three different, and diverse foundation models for vision tasks, which differ in the number of modalities for training (i.e., vision vs vision-language), strategy used for training (i.e., contrastive vs self-supervised learning), and backbones used (i.e., vision transformers vs convolutional neural networks). More conctretely, the DINO [7, 33] models employed in this work are ViT-based models trained in a self-supervised manner. CLIP [38] is a vision-language model, which trains a vision and a text encoder in a contrastive fashion. Last, VICReg [3] are CNNs trained in a self-supervised manner using contrastive learning.
Implementation details.
All linear probing heads are trained with a cross-entropy loss with the Adam [21] optimizer and a learning rate of $10^{-4}$, whose patience is set to $10$. The implementation of the training is based on PyTorch [34], and the conformal predictions methods on TorchCP [54]. For RAPS, we used $k_{\text{RAPS}}=2$ and $\lambda_{\text{RAPS}}=0.1$ as hyperparameters. Across all experiments, we use $\alpha=0.1$, except for CIFAR-10 for which we use $\alpha=0.05$.
C Relationship Between Model Performance and CP Metrics
Table 4, Table 5, and Table 6 report the numerical values corresponding to Fig. 1, showcasing the relationship between the linear probing accuracy of each model, and the CP performance in terms of set size and coverage gap for ImageNet, CIFAR-10, and CIFAR-100, respectively.
D Model performance for the optimal temperature
Figures 8 and 9 showcase the impact of temperature $T$ on the ECE, set size, minimum class-conditional coverage, and coverage gap across all foundation models. The observations on these figures strongly align with the findings on the main paper as, regardless of the model, APS yields optimal coverage gap close to the optimal temperature point. Indeed, looking at the different models (i.e., DINO-based, VICReg and CLIP), we can observe that the behaviour of RAPS (in terms of coverage gap), strongly varies across models. Additionally, Fig. 10 shows the evolution of $q_{\alpha}$, the threshold on the conformal scores $s$ before and after calibration. When applying temperature scaling, the distribution of softmax across classes approaches a uniform distribution, lowering the scores for the most likely classes. Intuitively, this means that more classes will need to be included in the set to ensure coverage, leading to a decrease in threhsold $q_{\alpha}$. This decrease is consistent with the observed increase in set size, which can be seen in Fig. 11, showing the difference in set size when applying temperature scaling for APS on CLIP. Note that this trend holds across models, as a similar behaviour for DINO is observed in Fig. 7 in the main paper. Note that both Figures 7 and 11 only show cases where the set sizes are different between the calibrated and uncalibrated model.
Furthermore, we report in Table 7 a more complete version of the results presented in Table 2, containing all explored models.
E Domain Shift
Table 8 reports the numerical values for Fig. 4, showcasing APS’s strong performance in terms of marginal and conditional coverage under distribution shift, at the cost of decreasing the set efficiency. Fig. 12, Fig. 13, and Fig. 14 depict the distribution of class-conditional coverage for DINOv2-B, VICReg (ResNet 50x2), and CLIP, respectively. The kernel density estimate plots show the class-conditional coverage for a distribution shift with ImageNet-A (top), ImageNet-R (middle), and ImageNet-V2 (bottom), for APS (left) and RAPS (right). These results clearly demonstrate the stronger resistance to distribution shift for APS compared to RAPS, even when both methods show a minimum class-conditional coverage of 0.
F Few-Shot Adaptation
We present in Table 9 a detailed version of Table 3, where the OOD section (in Table 3) corresponds to the average across all ImageNet variants.
AvgSize ($\downarrow$) MCCC ($\uparrow$)
Accuracy ($\uparrow$) LAC APS RAPS LAC APS RAPS
DINO-S 0.7492 3.27 10.02 4.23 0.072 0.059 0.074
DINO-B 0.7743 2.58 10.73 3.03 0.072 0.058 0.073
DINOv2-S 0.7948 1.87 6.66 2.12 0.073 0.057 0.071
DINOv2-B 0.8393 1.36 7.50 1.67 0.076 0.057 0.065
DINOv2-L 0.8624 1.21 6.77 1.50 0.079 0.057 0.063
DINOv2-G 0.8666 1.18 4.08 1.44 0.079 0.058 0.063
VICReg (ResNet-50) 0.7207 4.24 14.22 5.73 0.070 0.058 0.071
VICReg (ResNet-50x2) 0.7570 3.08 13.49 3.89 0.070 0.056 0.070
VICReg (ResNet-200x2) 0.7804 2.50 12.05 2.98 0.071 0.056 0.071
CLIP 0.7201 3.03 9.50 3.73 0.069 0.058 0.070
Table 4: Linear probing accuracy, average set size, and minimum class-conditional coverage for all models on ImageNet.
AvgSize ($\downarrow$) MCCC ($\uparrow$)
Accuracy ($\uparrow$) LAC APS RAPS LAC APS RAPS
DINO-S 0.8711 1.36 1.68 1.53 0.023 0.010 0.017
DINO-B 0.9165 1.14 1.41 1.30 0.024 0.009 0.013
DINOv2-S 0.9140 1.13 1.41 1.29 0.024 0.009 0.011
DINOv2-B 0.9511 1.01 1.22 1.14 0.025 0.008 0.009
DINOv2-L 0.9848 0.95 1.03 1.01 0.020 0.008 0.009
DINOv2-G 0.9932 0.95 0.99 0.98 0.020 0.008 0.008
VICReg (ResNet-50) 0.8506 1.49 1.84 1.69 0.021 0.011 0.018
VICReg (ResNet-50x2) 0.8749 1.32 1.65 1.50 0.022 0.010 0.014
VICReg (ResNet-200x2) 0.8481 1.47 1.85 1.66 0.022 0.011 0.017
CLIP 0.8823 1.33 1.71 1.51 0.021 0.009 0.013
Table 5: Linear probing accuracy, average set size, and minimum class-conditional coverage for all models on CIFAR-10.
AvgSize ($\downarrow$) MCCC ($\uparrow$)
Accuracy ($\uparrow$) LAC APS RAPS LAC APS RAPS
DINO-S 0.6668 3.46 5.64 4.29 0.050 0.037 0.055
DINO-B 0.7379 2.19 4.07 2.50 0.050 0.034 0.050
DINOv2-S 0.7515 2.06 4.06 2.40 0.049 0.035 0.050
DINOv2-B 0.8116 1.39 3.40 1.76 0.054 0.034 0.042
DINOv2-L 0.8885 1.02 2.37 1.35 0.061 0.035 0.036
DINOv2-G 0.9235 0.95 1.93 1.21 0.067 0.035 0.036
VICReg (ResNet-50) 0.6503 3.83 6.15 4.83 0.050 0.034 0.056
VICReg (ResNet-50x2) 0.6902 3.06 5.30 3.78 0.054 0.035 0.056
VICReg (ResNet-200x2) 0.6461 3.95 6.23 4.91 0.053 0.037 0.058
CLIP 0.6667 3.52 6.28 4.18 0.050 0.037 0.055
Table 6: Linear probing accuracy, average set size, and minimum class-conditional coverage for all models on CIFAR-100.
(a) DINO-S
(b) DINO-B
(c) DINOv2-S
(d) DINOv2-B
(e) DINOv2-L
(f) DINOv2-G
Figure 8: Impact of the temperature $T$ on the ECE, set size, MCCC, and CovGap, for DINO-S, DINO-B, DINOv2-S, DINOv2-B, DINOv2-L, and DINOv2-G on ImageNet. Red vertical line indicates the uncalibrated model performance ($T=1$). Green vertical line indicates the model performance for the optimal temperature.
(a) VICReg (ResNet 50)
(b) VICReg (ResNet 50x2)
(c) VICReg (ResNet 200x2)
(d) CLIP
Figure 9: Impact of the temperature $T$ on the ECE, set size, MCCC, and CovGap, for VICReg (ResNet 50), VICReg (ResNet 50x2), VICReg (ResNet 200x2), and CLIP on ImageNet. Red vertical line indicates the uncalibrated model performance ($T=1$). Green vertical line indicates the model performance for the optimal temperature.
(a) LAC
(b) APS
(c) RAPS
Figure 10: ECE and average $q_{\alpha}$ threshold before and after calibration for (a) LAC, (b) APS, and (c) RAPS on ImageNet. Each pair represents one model before (square) and after (circle) calibration.
Figure 11: Difference in conformal set size (i.e., efficiency) when applying temperature scaling on APS and CLIP ($T=1.1$).
ECE ($\downarrow$) AvgSize ($\downarrow$) MCCC ($\uparrow$)
(×$\times 10^{-2}$) LAC APS RAPS LAC APS RAPS
DINO-S 5.59/3.02 3.27/3.26 10.02/11.78 4.23/3.99 0.433/0.427 0.477/0.480 0.412/0.415
DINO-B 3.94/2.12 2.58/2.55 10.73/13.14 3.03/3.03 0.416/0.398 0.521/0.517 0.391/0.409
DINOv2-S 2.71/1.37 1.87/1.87 6.66/8.34 2.12/2.19 0.370/0.351 0.520/0.538 0.384/0.409
DINOv2-B 2.85/1.81 1.36/1.36 7.50/10.46 1.67/1.77 0.414/0.406 0.476/0.482 0.489/0.498
DINOv2-L 2.99/1.81 1.21/1.21 6.77/9.67 1.50/1.59 0.326/0.336 0.502/0.522 0.453/0.487
DINOv2-G 3.66/2.10 1.18/1.18 4.08/5.69 1.44/1.50 0.244/0.243 0.458/0.484 0.350/0.375
VICReg (RN-50) 3.12/1.90 4.24/4.19 14.22/16.6 5.73/5.49 0.472/0.468 0.566/0.570 0.453/0.456
VICReg (RN-50x2) 2.34/2.16 3.08/3.09 13.49/16.06 3.89/3.80 0.442/0.430 0.549/0.557 0.445/0.446
VICReg (RN-200x2) 2.21/1.90 2.50/2.49 12.05/14.61 2.98/3.01 0.440/0.430 0.567/0.566 0.462/0.471
CLIP 2.70/1.63 3.03/3.01 9.50/11.11 3.73/3.53 0.434/0.441 0.556/0.564 0.418/0.414
Table 7: Quantitative impact of calibrating vision foundation models. Average size and minimum class conditional coverage are reported between uncalibrated/calibrated models ($T=1$ and $T=1.1$ respectively) for several models, with ImageNet as benchmark dataset.
Set size ($\downarrow$) Coverage CovGap ($\downarrow$)
LAC APS RAPS LAC APS RAPS LAC APS RAPS
ImageNet DINOv2-B 1.36 7.50 1.67 0.900 0.900 0.900 0.414 0.476 0.489
VICReg (RN 50x2) 3.08 13.49 3.89 0.900 0.900 0.900 0.442 0.549 0.445
CLIP 3.03 9.50 3.73 0.900 0.901 0.900 0.434 0.556 0.418
ImageNet-V2 DINOv2-B 1.38 9.47 1.74 0.879 0.892 0.881 0.081 0.187 0.068
VICReg (RN 50x2) 3.15 12.54 3.89 0.867 0.871 0.870 0.130 0.175 0.143
CLIP 2.88 6.79 3.58 0.877 0.864 0.882 0.023 0.034 0.029
ImageNet-Sketch DINOv2-B 1.35 10.90 1.91 0.884 0.904 0.893 0.221 0.322 0.275
VICReg (RN 50x2) 3.49 21.98 4.51 0.826 0.879 0.830 0.119 0.223 0.114
CLIP 3.16 12.98 3.99 0.877 0.900 0.878 0.182 0.293 0.185
ImageNet-A DINOv2-B 1.63 25.30 3.55 0.646 0.897 0.734 0.000 0.000 0.000
VICReg (RN 50x2) 8.43 59.83 7.82 0.531 0.879 0.511 0.000 0.000 0.000
CLIP 6.79 39.42 6.45 0.648 0.912 0.624 0.000 0.000 0.000
ImageNet-R DINOv2-B 1.31 7.49 1.94 0.867 0.910 0.879 0.282 0.532 0.408
VICReg (RN 50x2) 5.84 27.36 5.65 0.761 0.897 0.742 0.198 0.429 0.186
CLIP 3.38 11.93 4.13 0.893 0.930 0.890 0.422 0.604 0.386
Table 8: Evaluation under domain-shift. Set size($\downarrow$), coverage($\uparrow$), and MCCC($\uparrow$) across three CP methods and three foundation models.
(a) DINOv2-B, ImageNet-A
(b) DINOv2-B, ImageNet-R
(c) DINOv2-B, ImageNet-Sketch
Figure 12: Domain shift analysis. Distribution of class-conditional coverages for DINOv2-B on ImageNet-A, ImageNet-R, and ImageNet-Sketch: APS (left) and RAPS (right).
(a) VICReg (RN 50x2), ImageNet-A
(b) VICReg (RN 50x2), ImageNet-R
(c) VICReg (RN 50x2), ImageNet-Sketch
Figure 13: Domain shift analysis. Distribution of class-conditional coverages for VICReg (RN 50x2) on ImageNet-A, ImageNet-R, and ImageNet-Sketch: APS (left) and RAPS (right).
(a) CLIP, ImageNet-A
(b) CLIP, ImageNet-R
(c) CLIP, ImageNet-Sketch
Figure 14: Domain shift analysis. Distribution of class-conditional coverages for CLIP on ImageNet-A, ImageNet-R, and ImageNet-Sketch: APS (left) and RAPS (right).
Set size ($\downarrow$) CovGap ($\downarrow$)
LAC APS RAPS LAC APS RAPS
ImageNet ZS 2.81 10.07 3.23 0.086 0.069 0.084
ZSLP 2.17 6.60 2.45 0.076 0.060 0.074
CLAP [45] 2.15 6.83 2.45 0.077 0.060 0.074
CoOp [58] 2.44 6.32 2.79 0.080 0.066 0.080
KgCoOp [14] 2.47 7.18 2.84 0.081 0.065 0.081
-V2 ZS 4.64 17.34 5.72 0.128 0.124 0.128
ZSLP 3.89 12.51 4.81 0.126 0.121 0.125
CLAP [45] 3.72 12.72 4.62 0.126 0.122 0.125
CoOp [58] 4.08 11.26 4.87 0.126 0.123 0.126
KgCoOp [14] 4.09 12.65 4.93 0.126 0.123 0.126
-Sketch ZS 14.74 37.18 20.39 0.099 0.090 0.099
ZSLP 16.46 37.27 22.79 0.098 0.091 0.098
CLAP [45] 15.20 34.90 21.29 0.098 0.090 0.098
CoOp [58] 16.06 34.36 22.12 0.098 0.091 0.099
KgCoOp [14] 15.46 34.35 21.74 0.098 0.092 0.099
-A ZS 9.42 16.59 11.28 0.094 0.085 0.095
ZSLP 10.92 18.63 12.47 0.087 0.080 0.088
CLAP [45] 9.36 16.56 11.38 0.087 0.082 0.087
CoOp [58] 9.23 14.65 10.72 0.087 0.084 0.090
KgCoOp [14] 8.66 14.39 10.19 0.093 0.084 0.096
-R ZS 1.92 5.77 2.31 0.059 0.039 0.054
ZSLP 2.07 5.62 2.36 0.058 0.041 0.056
CLAP [45] 1.92 5.46 2.26 0.059 0.039 0.053
CoOp [58] 2.06 5.11 2.40 0.061 0.042 0.063
KgCoOp [14] 1.96 5.32 2.29 0.060 0.041 0.058
Table 9: Set size and CovGap for CLIP (ViT-B backbone) on ImageNet and its variants.