### v2

# Are foundation models for computer vision good conformal predictors?

Abstract
Recent advances in self-supervision and contrastive learning have brought the performance of foundation models to unprecedented levels in a variety of tasks. Fueled by this progress, these models are becoming the prevailing approach for a wide array of real-world vision problems, including risk-sensitive and high-stakes applications. However, ensuring safe deployment in these scenarios requires a more comprehensive understanding of their uncertainty modeling capabilities, which has been barely explored. In this work, we delve into the behaviour of vision and vision-language foundation models under Conformal Prediction (CP), a statistical framework that provides theoretical guarantees of marginal coverage of the true class. Across extensive experiments including popular vision classification benchmarks, well-known foundation vision models, and three CP methods, our findings reveal that foundation models are well-suited for conformalization procedures, particularly those integrating Vision Transformers. We also show that calibrating the confidence predictions of these models, a popular strategy to improve their uncertainty quantification, actually leads to efficiency degradation of the conformal set on adaptive CP methods. Furthermore, few-shot adaptation of Vision-Language Models (VLMs) to downstream tasks, whose popularity is surging, enhances conformal scores compared to zero-shot predictions. Last, our empirical study exposes APS as particularly promising in the context of vision foundation models, as it does not violate the marginal coverage guarantees across multiple challenging, yet realistic scenarios.
1 Introduction
Large-scale pre-trained vision foundation models, such as DINOv2 [40], as well as those integrating text, such as CLIP [46], are driving a new learning paradigm in machine learning, achieving unprecedented results on a broad spectrum of tasks. Despite their desirable zero-shot and generalization capabilities, recent evidence has pointed out to the existence of bias and factual errors in these models [58], which transcend the field of computer vision [52, 4]. For example, the original CLIP paper [46] demonstrated gender and race biases in certain zero-shot tasks, whereas [58] identified that CLIP-based models are not always better calibrated than other arguably simpler ImageNet models. Furthermore, [36] recently showed that adapted models magnify the miscalibration issue compared to the zero-shot setting, yielding overconfident predictions. These problems underscore widespread societal concerns surrounding the reliable deployment and use of foundation models in sensitive contexts, such as decision-making processes in critical scenarios, e.g., healthcare or security applications.
A popular solution to quantify the uncertainty present in the predictions of deep models is calibration. In this setting, the proposed strategies aim at reducing the discrepancies between model predictions and the actual correctness probability. Temperature Scaling (TS) [16], a simple variant of Platts Scaling [44], provides a simple post-processing approach to adjust the softmax probability scores of the trained models. Other lines of methods have proposed training objectives to enforce the model to produce less confident scores, either in the predictions space [45, 43], logits [32, 31], or modifying the ground truth labels [34, 35].
Conformal Prediction (CP) [60], an alternative strategy to quantify the uncertainty [49, 48, 1], is a statistical framework which offers several advantages over calibration methods. First, unlike most methods in calibration, CP works directly on the model predictions, offering an appealing solution for black-box models. Second, instead of simply modeling the correctness of the predicted probability, CP produces a set of predictions, including the most likely classes, which can be of much interest in certain problems. Last, CP methods have theoretical guarantees for the marginal coverage of the true class within the predicted set, under several assumptions, contrary to calibration approaches.
Due to these properties, CP is gaining attention to conformalize the predictions of deep models [2, 1, 11, 50]. Nevertheless, albeit the efforts to study CP in large language foundation models [15], its impact on vision foundation models has been unexplored, besides the significant implications it may have on a variety of vision problems. The aim of this study is to shed light and provide important insights into this direction. To achieve this, we conducted an extensive empirical analysis of the performance of three common CP methods on eleven popular vision foundation models across multiple vision datasets. Our extensive experiments further explore common situations encountered in practice, assessing the impact on CP: under distributional drifts, after confidence calibration and in few-shot adaptation to novel downstream tasks. Our key observations are:
  1. (i)
    Vision, and vision-language foundation models seem to better conformalize compared to their more traditionally (fully-supervised) trained counterparts.
  2. (ii)
    Across different foundation models, those including visual transformers, such as DINO and CLIP, lead to better conformal results compared to models integrating convolutional neural networks.
  3. (iii)
    Across all the experiments, Adaptive Prediction Sets (APS) is the best CP approach in terms of empirical coverage, while Regularized Adaptive Prediction Sets (RAPS) presents the best alternative from a conformal set size standpoint.
  4. (iv)
    Under distributional shifts, APS exhibits the highest robustness among CP methods in terms of coverage guarantes, albeit decreasing its set efficiency.
  5. (v)
    Confidence calibration decreases the efficiency of conformal sets, but typically improves coverage gap.
  6. (vi)
    Few-shot adaptation of vision-language models (VLMs) yields better conformalization than zero-shot predictions in ID data, with marginal gains on OOD.
2 Related Work
Foundation models for computer vision.  The landscape of foundation models has rapidly evolved in the last years. Traditionally, pre-trained convolutional networks based on ResNet architectures [18] were the main models used by the community. However, driven by the unprecedented advances in language models, e.g., GPT [7] or LLaMA [57], as well as the vast availability of image data online, there are groundbreaking advances in unimodal [8, 40, 26] and multimodal [46] foundation models for vision tasks, commonly based on vision transformers. These large pre-trained models aim to generalize across a broad span of visual tasks by pre-training on massive, diverse image datasets, exhibiting strong zero-shot and generalization capabilities to new tasks. For example, vision foundation models such as DINO [8, 40] rely on self-supervised learning strategies on large datasets, leading to excellent semantic understanding of visual content. On the other hand, CLIP [46] bridges the gap between language and vision modalities through contrastive learning, effectively allowing the model to understand images in the context of natural language prompts and enabling zero-shot capabilities.
Quantifying the uncertainty of the predictions of deep networks has recently garnered considerable interest. From a calibration standpoint, popular strategies include post-hoc approaches [16, 12, 24], which map the logits or softmax predictions to smoother distributions, and explicit learning objectives [31, 32, 35, 34, 19, 37], which are integrated into the loss function. Nevertheless, a main limitation of calibration methods is that they lack theoretical guarantees of model performance. In contrast, CP has recently emerged as a promising alternative, which provides marginal coverage guarantees over unseen test samples [61, 51]. Specifically, CP resorts to a non-conformity score function (i.e., a measure of how “different” a particular data point is compared to a CP calibration dataset) to produce a finite prediction set, which is guaranteed to contain the true label with a user-specified confidence level. A central objective of the CP literature has been to improve either the set efficiency (i.e., smaller set sizes) or the class conditional coverage. For this purpose, several non-conformity scores have been presented [56, 48, 1, 49, 13, 11, 55], with [48, 1, 49] being popular methods widely studied. A straightforward solution directly uses the raw class softmax predictions to generate the prediction sets [49]. Adaptive Prediction Sets (APS) [48] provides an adaptive version, computing non-conformity scores by accumulating sorted softmax probabilities in descending order. To further improve the efficiency, RAPS [1] introduces an explicit regularization term, which penalizes non-conformity scores for unlikely classes.
However, a main limitation of existing evaluations is the focus on more traditional models, usually trained on data collection that falls within the calibration and test data points distribution. Despite this transfer learning framework not necessarily affecting the marginal guarantees provided by CP, how it affects its efficiency and conditional coverage remains to be explored. Thus, quantifying the uncertainty of their predictions is paramount given the rising popularity of foundation models in strategic domains. However, whereas uncertainty quantification from a calibration perspective has been scarcely studied [36, 66, 59], its exploration under CP is, to our knowledge, overlooked.
3 Background
3.1 Conformal Prediction Framework
Let $\mathcal{X}$ and $\mathcal{Y}$ denote the input and output space, respectively. We assume access to a calibration set $\mathcal{D}_{cal}=\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{n}$ of $n$ independent and identically distributed (i.i.d.) samples, where each $\mathbf{x}_{i}=(p_{ik})_{1\leq k\leq K}$ represents the black-box probabilities and $y_{i}\in\mathcal{Y}=\{1,2,...,K\}$ is the associated label. The goal of CP is to construct a prediction set $\hat{C}(\mathbf{x}_{n+1})\subseteq\mathcal{Y}$ for a new test input $\mathbf{x}_{n+1}$ such that it contains the true label $y_{n+1}$ with a user pre-specified coverage probability $1-\alpha$, where $\alpha\in(0,1)$ denotes the error level.
The core idea of CP is to assess the degree to which a new sample conforms to the underlying distribution of the calibration data by computing non-conformity scores. Let $S(\mathbf{x},y)$ be a non-conformity measure (or scoring function) that assigns a score $s_{i}$ to each $(\mathbf{x}_{i},y)$ pair, quantifying how unusual the pair is relative to the rest of the data. Given the non-conformity scores for all calibration examples and a new input $\mathbf{x}_{n+1}$ (unseen in the calibration set), the conformal prediction set $C(\mathbf{x}_{n+1})$ is defined as:
$C(\mathbf{x}_{n+1})=\left\{y\in\mathcal{Y}:S(\mathbf{x}_{n+1},y)\leq q_{\alpha}\right\},$ (1)
where $q_{\alpha}$ is the $1-\alpha$ quantile of the non-conformity scores on the calibration set, obtained with the observed labels:
$q_{\alpha}=\textsc{Quantile}\Big{(}\{S(\mathbf{x}_{i},y_{i})\}_{i=1}^{N},\frac{\lceil(n+1)(1-\alpha)\rceil}{n}\Big{)}$ (2)
Coverage Guarantees.  A key property of conformal prediction is its finite-sample coverage guarantee. This property ensures that the prediction sets achieve the desired coverage probability marginally over $\mathcal{X}$ and $\mathcal{Y}$, irrespective of the underlying data distribution, as long as the calibration and test data are exchangeable [61]. Formally, for any $1-\alpha$, conformal predictors satisfy:
$\mathbb{P}(y_{n+1}\in C(\mathbf{x}_{n+1}))\geq 1-\alpha.$ (3)
This property is crucial for applications requiring reliable uncertainty quantification, particularly where distributional assumptions (e.g., Gaussianity) may not hold.
Tightness of Prediction Sets.  Conformal prediction guarantees valid marginal coverage. Nevertheless, the efficiency of the prediction sets, i.e., their size, depends on the choice of the non-conformity measure. Choosing an appropriate $S(\mathbf{x},y)$ is key to balancing the trade-off between coverage and tightness of the prediction sets. In practice, we aim to minimize the size of the prediction sets while maintaining the desired coverage probability.
3.2 Non-conformity scores
Least Ambiguous Classifier (LAC) [49]  aims to construct the smallest possible set under the assumption that the output is correct. Intuitively, it can be interpreted as a thresholding of the output probabilities for each category. Thus, the non-conformity score can be constructed as:
$\mathcal{S}_{\text{\tiny{LAC}}}(\mathbf{x},y)=1-x_{k=y}.$ (4)
LAC also provides notable efficiency in scenarios using an imperfect classifier. However, it lacks adaptability, e.g., in under-represented categories or uncertain predictions.
Adaptive Prediction Sets.  APS [48] provides a non-conformity score that leverages the accumulated confidence in the ordered probability predictions. Thus, APS is known to be an adaptive score, whose main objective is enhancing the coverage of uncertain predictions by sacrificing efficiency. Formally, APS is expressed as:
$\mathcal{S}_{\text{\tiny{APS}}}(\mathbf{x},y)=\rho(\mathbf{x},y)+x_{k=y}\cdot u,$ (5)
where $\rho(\mathbf{x},y)$ is the accumulated confidence of the categories more likely than the evaluated label $y$, i.e., $\rho(\mathbf{x},y)=\sum_{k^{\prime}\in\mathcal{Y^{\prime}}(\mathbf{x},y)}x_{k=k^{\prime}}$, with $\mathcal{Y^{\prime}}(\mathbf{x},y)=\{k|x_{k}>x_{k=y}\}$. Adaptive methods usually include $u\in\{0,1\}$, as a random variable to break ties to archieve exact marginal coverage.
Regularized Adaptive Prediction Sets.  RAPS [1] builds upon APS by adding a regularization term to enforce smaller predicted sets. Thus, APS score is modified to penalize the confidence of introducing additional, unlikely categories, after a certain set size is met:
$\mathcal{S}_{\text{\tiny{RAPS}}}(\mathbf{x},y)=\rho(\mathbf{x},y)+x_{k=y}\cdot u+\lambda(o(\mathbf{x},y)-k_{reg})^{+}$ (6)
where $\lambda,k_{reg}\geq 0$ are hyper-parameters controlling the penalty strength, $o_{x}(y)$ is the rank of the sorted label, $o(\mathbf{x},y)=|\mathcal{Y^{\prime}}(\mathbf{x},y)|+1$, and $(\cdot)^{+}$ denotes the positive part.
We refer the reader to the different works [1, 49, 48] for the respective conformal calibration coverage guarantees.
4 Experiments
4.1 Experimental Setup
Models: We employ a total of 11 foundation models: two DINO [8] (DINO-S and DINO-B), four DINOv2 [40] (DINOv2-S, DINOv2-B, DINOv2-L, and DINOv2-G), three VICReg [3] (with ResNet-50, ResNet-50x2, and ResNet-200x2), and two VLMs (CLIP [46] and MetaCLIP [65]) based on ViT-B. Our main analysis is conducted on three popular vision datasets: CIFAR-10 [28], CIFAR-100 [28], and Imagenet [10], including its versions integrating domain shifts [21, 22, 62, 47]. Each dataset is split into two sets: one for training, and one for the conformal experiments. The latter is then split into one calibration set to tune the CP method, and one test set for evaluation. For few-shot, we adhere to the emerging CLIP few-shot literature [68, 53, 69], and evaluate models on 10 additional fine-grained and general concepts classification benchmarks: SUN397 [64], FGVCAircraft [33], EuroSAT [20], StanfordCars [27], Food101 [6], OxfordPets [41], Flowers102 [39], Caltech101 [14], DTD [9], and UCF101 [54].
Metrics: We resort to conformal set size (a.k.a. efficiency), minimum class-conditional coverage (MCCC), coverage gap (CovGap) and marginal coverage to benchmark the different methods. More details about metrics and implementation details can be found in Appendix A and Appendix B.
Figure 1: Relationship between the linear probing model accuracy and conformal set size (top) and the coverage gap (bottom) across different tasks of increasing complexity. From left to right: CIFAR-10, CIFAR-100 and ImageNet.
Adaptation to target tasks.  The foundation models used in this study have been pre-trained using different strategies. Nevertheless, they need to be adapted for novel tasks, as their pre-trained versions do not accommodate classification tasks, i.e., there is no classification head. To do this, foundation models are frozen, and a linear probing (LP) head (one linear layer followed by a softmax activation function) is trained on each dataset by optimizing a cross-entropy loss (more details in Appendix B).
4.2 Results
To gain insights into the factors influencing the efficacy of CP in vision foundation models, we design four experiments. First, we explore the impact of CP in standard scenarios, where a large calibration set is available to conform to the predictions of different models. Then, we challenge the status quo of CP and alter the conformal sets to accommodate real-world scenarios by including domain shifts. Furthermore, since confidence calibration is significantly linked to CP, we explore the impact of model calibration on CP performance. Last, we examine CP when adapting a very popular VLM, i.e., CLIP, to novel tasks.
4.2.1 Performance in the General Setting
First, we study the performance of 11 vision foundation models paired with CP methods under the standard setting, and on the three datasets, which presents ideal conditions: a sufficiently large calibration set, and absence of distributional drifts between calibration and test sets. We aim to determine whether we can prescribe a winner solution in this scenario and which factors can help identifying it.
Fig. 1 depicts the relationship between the linear probing performance of each model and the conformal set size (top), and coverage gap (bottom). The initial observation highlights a clear trend: higher-performance models tend to produce smaller prediction sets, regardless of the conformal method used. Nevertheless, while set efficiency (i.e. size) is typically considered as a sufficient condition in most prior literature in conformal prediction [1], our analysis revealed that the relation with the accuracy does not consistently hold when examining other metrics. In particular, Fig. 1 (bottom) exposes that different CP methods yield mixed results for the coverage gap, which do not correlate directly with accuracy across all CP approaches. While APS appears to be almost unaffected by the model performance, RAPS clearly benefits from a strong performance, and LAC is negatively affected by more accurate models as the dataset becomes more complex.
Regarding comparison between the CP methods, if we consider the set size, APS is clearly outperformed by the other approaches, particularly on ImageNet, whereas LAC provides the smallest prediction sizes, closely followed by RAPS. Indeed, RAPS is specifically designed to reduce the conformal set size of APS. However, the coverage gap results indicate that this comes at the cost of increasing the range of the class-conditional coverage. Below, we analyze the underlying causes that may explain this behaviour.
RAPS class-conditional coverage, and therefore coverage gap, are more sensitive to the model’s accuracy. Let us assume we have two models, $M_{1}$ and $M_{2}$, in a multi-class classification problem, whose accuracies are $Acc_{1}$ and $Acc_{2}$, respectively. For each class $y\in\mathcal{Y}$, we refer to $C_{M_{i}}(y)=\mathbb{P}(Y=y|Y\in\mathcal{S}_{M_{i}}(X))$ as the class-conditional coverage for $y$ under model $M_{i}$, which measures the probability that predictions include the true label when the true label is $y$. Furthermore, let $\delta_{M_{1}}=\min_{y\in\mathcal{Y}}C_{M_{1}}(y)$ and $\delta_{M_{2}}=\min_{y\in\mathcal{Y}}C_{M_{2}}(y)$ denote the minimum class-conditional coverage achieved by each model. Under this scenario, we argue that due to the penalty in RAPS, if $Acc_{1}<Acc_{2}$, then $\delta_{M_{1}}<\delta_{M_{2}}$ when using RAPS as a conformal prediction method. In particular, $M_{1}$ (with lower accuracy) needs to expand its prediction sets for certain classes to meet the marginal coverage target 1-$\alpha$. However, the penalty term encourages small prediction sets, limiting an excessive number of classes. Thus, for some difficult classes, model $M_{1}$ may still potentially fail to meet the target coverage, as the enforced penalty discourages overly large sets. This ultimately results in lower coverage rates for those specific classes than $M_{2}$. In contrast, since APS does not include any regularization term that encourages small set sizes, it will compensate for lower performing models by increasing its set sizes, ultimately attaining higher class-conditional coverages.
Figure 2: Comparison (APS vs RAPS) of the class-conditional coverage and set size for the class for which RAPS has the worst class-conditional coverage. Experiments performed on CIFAR-100. Models sorted (in ascending order) by their LP performance ($\min=0.65$ and $\max=0.92$), indicated by the size of the circles.
To confirm this hypothesis, we perform the following experiment. First, we identify the class with the lowest minimum class-conditional coverage obtained by RAPS for each model $m$, which we refer to as $y^{R}_{m}$, and find its corresponding set size (both represented blue circles in Fig. 2). Then, we identify the class-conditional coverage and set size of the class $y^{R}_{m}$ provided by APS across all models, whose values are shown as pink circles in both plots of Fig. 2. Last, we also include the minimum class-conditional coverage obtained by APS across each method, depicted with a cross. Note that the minimum class-conditional coverage from APS does not necessarily correspond to the coverage of class $y^{R}_{m}$. Upon close examination of these results, we observe that, indeed, for models presenting lower accuracy (those with smaller circles), the gap between the class-conditional coverage for the worst class in RAPS and the same class in APS is consistently larger than for more accurate models (larger circles). This empirical evidence shows that the minimum class-conditional coverage is substantially reduced because RAPS is constrained from expanding its predictions set size. This effect is more pronounced in less accurate models, where the true class may rank far away from the maximum allowable set size in the softmax predictions. In contrast, APS tolerates worse models by increasing the set size, which ultimately degrades the set efficiency but yields better class-conditional coverages.
Last, LAC presents structural differences with RAPS and APS, as it lacks an adaptive mechanism, relying on a uniform fixed threshold. Thus, LAC may yield inconsistent coverage rates across classes, resulting in high variability in the class-conditional coverage and thus in the coverage gap.
Set size ($\downarrow$) MCCC ($\uparrow$)
Acc ($\uparrow$) Thr APS RAPS Thr APS RAPS
$\text{ViT}_{\text{CLIP}}$ 72.01 3.03 9.50 3.73 0.434 0.556 0.418
$\text{ViT}_{\text{DINO-S}}$ 74.92 3.27 10.02 4.23 0.433 0.477 0.412
$\text{ViT}_{\text{MetaCLIP}}$ 75.80 2.39 9.43 2.84 0.479 0.535 0.467
$\text{ViT}_{\text{ImageNet}}$ 76.08 2.36 38.75 4.46 0.416 0.495 0.405
Table 1: SSL vs supervised learning. Results on ImageNet obtained by CLIP (ViT-B), MetaCLIP (ViT-B), and DINO-S (ViT-S) and a ViT-B trained in a supervised manner on ImageNet.
Following this analysis, we are also interested in determining whether a network pre-trained following a more traditional approach (i.e., standard supervised fine-tuning) offers similar conformal capabilities to self-supervised and contrastive ones. In particular, we select a ViT-B pre-trained on ImageNet, which is the same architecture as the visual encoder of the different foundation models. The results from this analysis (Table 1) reveal that, despite obtaining lower classification accuracy when using LP on the different foundation models, CP methods typically yield better performance than in $\text{ViT}_{\text{ImageNet}}$. These differences are significant under the APS approach, where the set size is significantly degraded on $\text{ViT}_{\text{ImageNet}}$. Moreover, the class-conditional coverage is also substantially affected, with nearly 6% decrease compared to the best model.
Figure 3: $\text{ViT}_{\text{ImageNet}}$ vs $\text{ViT}_{\text{CLIP}}$. Analyzing the difference in set size between a $\text{ViT}_{\text{CLIP}}$ and $\text{ViT}_{\text{ImageNet}}$. Equal set sizes not shown.
Figure 4: Evaluation under domain-shift. Set size($\downarrow$), coverage($\uparrow$), and MCCC($\uparrow$) across three CP methods and three foundation models. ImageNet versions are sorted based on OOD performance in [53].
To further delve into these differences, we compute, for each test sample, the difference between the conformal set size for APS when applied to $\text{ViT}_{\text{ImageNet}}$ and $\text{ViT}_{\text{CLIP}}$ models, whose distribution is depicted in Fig. 3 (additional results in Appendix E). These values confirm that set size differences are not derived from a small set of isolated outliers but from a considerably large group of samples that see their conformal set increase when using the ViT trained in a supervised manner. These results suggest that the strategies used to train foundation models yield better CP metrics, resulting in conformalized models that can be deployed more safely on critical scenarios. It is important to stress that this study is limited due to the different dataset scales used for training (i.e., ImageNet alone is insufficient to train a foundation model). Our goal, however, is to understand the conformalization properties of readily available pre-trained models, regardless of how they were pre-trained.
4.2.2 Impact under Distribution Shifts
The theoretical guarantees of the coverage for conformal prediction hold under the hypothesis that the calibration set and the test set are drawn from the same distribution, i.e., data exchangeability assumption. In this section, we analyze the impact of having calibration sets that present distributional shifts with regard to the testing set.
We resort to ImageNet and its different versions: ImageNet-R [21], ImageNet-A [22], ImageNet-Sketch [62] and ImageNet-V2 [47]. To introduce the distributional drift between the calibration and testing data, we adapt the pre-trained model to one of the ImageNet versions. Then, ImageNet is used as the calibration set to conform the model, which is later tested on the ImageNet version used for adaptation. This is repeated for each ImageNet variant.
APS [48] exhibits strong robustness against large distributional shifts, at the cost of substantially degrading efficiency. One would expect that adaptive CP methods, such as APS and RAPS, somehow mitigate domain shifts due to their adaptive nature. Nevertheless, Fig. 4 reveals several interesting observations, which contradict this intuition. First, we can observe that, when resorting to APS as CP method the coverage gap is consistently satisfied (or nearly satisfied) across all domain shifts and models (Fig. 4, middle). In contrast, RAPS generally shows very similar performances compared to LAC, obtaining lower marginal coverage under several models and domains, and substantially lower than APS. To understand this phenomenon, we now study how set sizes evolve across domains for the different methods (Fig. 4, left). We can easily observe that (i) APS yield the largest conformal sets across ImageNet domains, regardless of the model, and (ii) APS experiences the largest set increases when the complexity of the domain grows. Thus, as exposed in the previous Section, APS satisfied marginal coverage by substantially including more predicted classes, therefore increasing conformal set sizes.
Figure 5: Domain shift analysis. Distribution of class-conditional coverages for CLIP on ImageNet-A: APS (left) and RAPS (right).
ECE ($\downarrow$) AvgSize ($\downarrow$) MCCC ($\uparrow$)
(×$\times 10^{-2}$) LAC APS RAPS LAC APS RAPS
DINOv2-S 2.71/1.37 1.87/1.87 6.66/8.34 2.12/2.19 0.370/0.351 0.520/0.538 0.384/0.409
DINOv2-B 2.85/1.81 1.36/1.36 7.50/10.46 1.67/1.77 0.414/0.406 0.476/0.482 0.489/0.498
DINOv2-L 2.99/1.81 1.21/1.21 6.77/9.67 1.50/1.59 0.326/0.336 0.502/0.522 0.453/0.487
DINOv2-G 3.66/2.10 1.18/1.18 4.08/5.69 1.44/1.50 0.244/0.243 0.458/0.484 0.350/0.375
VICReg (RN-50x2) 2.34/2.16 3.08/3.09 13.49/16.06 3.89/3.80 0.442/0.430 0.549/0.557 0.445/0.446
VICReg (RN-100x2) 2.21/1.90 2.50/2.49 12.05/14.61 2.98/3.01 0.440/0.430 0.567/0.566 0.462/0.471
CLIP 2.70/1.63 3.03/3.01 9.50/11.11 3.73/3.53 0.434/0.441 0.556/0.564 0.418/0.414
MetaCLIP 2.63/1.99 2.39/2.40 9.43/11.31 2.84/2.84 0.479/0.477 0.535/0.541 0.467/0.469
Table 2: Quantitative impact of calibrating vision foundation models. Average size and minimum class conditional coverage are reported between uncalibrated/calibrated models ($T=1$ and $T=1.1$ respectively) for several models, with ImageNet as benchmark dataset. Arrows indicate a decrease () or an increase() in metric after calibration, color indicates better (green) and worse (red) performance.
Figure 6: Impact of the temperature $T$ on ECE ($\downarrow$), set size ($\downarrow$), MCCC ($\uparrow$), and CovGap ($\downarrow$) (CLIP on ImageNet). Green region indicates the area with ECE smaller than for $T=1$. Green dotted line indicates value of optimal calibration ($T=1.1$). More plots in Appendix D.
APS and RAPS are adaptive methods that produce similar minimum class-conditional coverage, as exhibited in ImageNet-A (Fig. 4, right). However, they showcase significant differences in the average coverage gap. In the following, we explore this phenomenon in detail. More concretely, we depict in Fig. 5 the distribution of the conditional class coverage values obtained by APS and RAPS on ImageNet-A for CLIP (more datasets and models in Appendix F). Interestingly, while both approaches see their minimum class-conditional coverage decrease, their distributions are completely different. Indeed, APS distributions exhibit a Gaussian shape, with a decreasing number of categories presenting lower conditional coverage as they separate from 1-$\alpha$. In contrast, the distribution in RAPS exposes a significantly worse scenario, where class-conditional coverages are almost uniformly spread, with a non-negligible amount of classes below the expected coverage, i.e., $1-\alpha$.
Another interesting and valuable observation is that the low performance of all CP methods is magnified when coupled with VICReg (i.e., larger set sizes and lower coverage), suggesting that this family of foundation models may suffer more under distribution shifts.
4.2.3 Does calibration affect Conformal Prediction?
Confidence calibration is a popular strategy to improve the uncertainty estimates of deep models. These techniques, which can be either added as a post-processing step [44, 16, 29] or integrated as a training regularizer [19, 5, 38], adapt the model softmax predictions to reflect their actual performance accurately. The exponentially growing adoption of vision and vision-language foundation models in critical areas makes integrating confidence calibration a natural progression, as evidenced by recent works [36, 66, 59]. Thus, in this section, we investigate this important issue, as the relationship between calibration and conformal prediction in vision foundation models remains largely unexplored. Specifically, we examine whether calibrating these models with the popular Temperature Scaling (TS) [16] affects the conformal performance of fixed and adaptive CP methods. Particularly, we apply TS to the ImageNet results in the general case (Section 4.2.1). As a proper validation set is not available, we evaluate the CP performance over a set of $T$ values (14 values from $0.85$ to $2$), and found that $T=1.1$ typically yielded well-calibrated models [^1] Note that our goal is not to obtain the best-calibrated model but evaluate the impact of calibration on CP. Furthermore, this value aligns with existing works [34, 24] using TS for similar datasets, e.g., TinyImageNet..
Our observations suggest that calibration decreases the efficiency of CP sets but increases the minimum class-conditional coverage, improving the coverage gap.
Confidence calibration typically smooths the distribution of the class softmax scores, which results in less confident predictions. Consequently, the dominant value in the predicted softmax vector is typically lower in calibrated models. Nevertheless, since CP methods provide theoretical guarantees (under the data exchangeability assumption [61]) to satisfy the target marginal coverage of $1-\alpha$, these changes in the softmax distributions affect the conformalization obtained pre-TS. We present in Table 2 the results for the average set size and minimum class-conditional coverage before and after scaling the logits with TS. Furthermore, we include the Expected Calibration Error (ECE) values to verify that model calibration has improved. These results show that if the model is calibrated, its prediction set size tends to be larger, particularly for adaptive CP approaches (RAPS, and more specifically, APS). Fig. 7 further delves into these results, where we plot the distribution of differences between the set size of samples before and after calibration (i.e., a point in the distribution is $\mathcal{C}(x_{i})-\mathcal{C}(x_{i}^{TS})$). We can identify that the overall larger size in APS is caused by a consistent efficiency degradation across samples and not a few atypical cases with large conformal sets. We also observe that class-conditional coverage is typically improved on calibrated models, particularly when they are conformalized by adaptive CP methods. Thus, these results suggest that calibrating vision foundation models decreases efficiency while marginally enhancing class-conditional coverage, particularly on adaptive CP.
Figure 7: Difference in conformal set size (i.e., efficiency) when applying temperature scaling on APS and DINOv2-B ($T=1.1$).
The best coverage gap obtained by APS [48] approximately aligns with the coverage gap achieved at the optimal calibration point, whereas RAPS [1] coverage gap strongly differs. Fig. 6 shows the impact of the temperature $T$ on the different CP metrics on CLIP conformalized predictions. An interesting observation is that at the optimal temperature value $T=1.1$, the CovGap of APS is very close to its optimal point [^2] Results on Appendix D show similar behavior for other models. (0.0567 vs 0.0561). In contrast, while the RAPS coverage gap at optimal calibration is 0.0701, it decreases to 0.0642 as $T$ increases. This suggests that we could anticipate the best coverage gap for APS by assessing the model calibration performance, which will likely be near optimal calibration. On the downside, APS conformal sets efficiency is degraded. As $T$ increases, the softmax distributions get smoother, explaining the degradation in efficiency, which is even more drastic for APS, whose set size monotonically increases with $T$. We concede that, whereas APS minimizes the coverage gap, it does so at the expense of increasing the conformal set size. However, we believe that increasing the size, up to some extent, while improving coverage gap is preferable in critical decision systems.
4.2.4 Effect on few-shot adapted models
Adapting zero-shot CLIP models for downstream tasks in a few-shot labeled regime is becoming increasingly popular in VLMs. These strategies can be mainly categorized into Prompt Learning [70, 17, 69], which optimize the set of text prompts given to the text encoder, and Adapters [67, 30, 23, 53, 70, 17], where only a limited set of learnable parameters atop embedding representations is updated. Thus, questioning whether adapting these models hinges the performance of CP methods is of paramount importance, as it addresses a foundational aspect of effectively integrating uncertainty quantification with modern architectures. To assess the impact of CLIP adaptation in conformal prediction, we resort to representative methods of the few-shot adaptation families presented above, and follow standard adaptation and evaluation protocols in the literature [53]: (i) for out-of distribution (OOD), we adapt CLIP on few-shot samples ($M=16$) from ImageNet, and evaluate on ImageNet and its variants, and (ii) for in distrbution (ID) we adapt on few-shot samples and evaluate on the validation set from the same dataset.
Set size ($\downarrow$) CovGap ($\downarrow$)
LAC APS RAPS LAC APS RAPS
ID ZS 5.29 6.98 6.49 0.114 0.094 0.102
ZSLP 2.09 3.43 2.53 0.087 0.060 0.064
CLAP [53] 2.13 3.52 2.57 0.088 0.060 0.065
CoOp [70] 2.07 2.87 2.47 0.083 0.059 0.067
KgCoOp [17] 2.10 3.11 2.53 0.085 0.059 0.065
OOD ZS 7.68 19.22 9.93 0.095 0.085 0.094
ZSLP 8.34 18.51 10.61 0.092 0.083 0.092
CLAP [53] 7.55 17.41 9.89 0.093 0.083 0.091
CoOp [70] 7.86 16.35 10.03 0.093 0.085 0.095
KgCoOp [17] 7.54 16.68 9.79 0.094 0.085 0.095
Table 3: Few-shot adaptation (16-shots). Set size and CovGap for CLIP (ViT-B backbone) on in-distribution (average over 11 datasets) and out-of-distribution (average over ImageNet versions). Results with 4-shots are provided in Appendix, Table 10.
Few-shot VLM adaptation render pre-trained models better conformalized in the ID scenario (Table 3). First, we observe that, across both ID and OOD scenarios, both Adapters and Prompt Learning yield smaller conformal set sizes and higher Coverage gap than zero-shot CLIP, regardless of the CP method, with Prompt Learning yielding slightly better performances. In contrast, on the OOD scenario, only APS consistently enhances the set efficiency of the ZS model, with all the few-shot adaptation methods obtaining scarce coverage gap improvements across CP methods. These observations are related to the findings from the previous section, which suggested that better-calibrated models lead to larger conformal sets. Indeed, recent evidence [36] demonstrated that few-shot CLIP adaptation methods deteriorate the confidence estimates compared to ZS predictions, which, following our observations, should result in smaller conformal sets, as validated in Table 3. Further details are provided in Appendix G.
5 Conclusion
In this study, we aimed at answering the question: Which CP method and model should I use, and what can I expect, in the era of vision foundation models? Our findings revealed that vision foundation models yield better conformal metrics than their traditional pre-trained counterparts, with models integrating visual transformers outperforming those that use convolutional neural networks, particularly under domain shifts. Furthermore, we observed that several common situations encountered in practice (i.e., presence of distributional drifts and models undergoing confidence calibration) are indeed detrimental for some CP approaches. Interestingly, an adaptive CP method, i.e., APS, exhibited stronger robustness to these scenarios, particularly in terms of conditional coverage, but at the expense of degrading the set efficiency.
Acknowledgments
This work has benefited from state financial aid, managed by the Agence Nationale de Recherche under the investment program integrated into France 2030, project reference ANR-21-RHUS-0003. This work was partially supported by the ANR Hagnodice ANR-21-CE45-0007. This work was partially supported by the ANR Hagnodice ANR-23-CE45-0029. This work was granted access to the HPC resources of IDRIS under the allocation 2023-AD011014802 made by GENCI. We gratefully acknowledge the DATAIA program for supporting JD as a visiting professor at Université Paris-Saclay.
Supplementary Material
A Metrics
Given a pre-specified coverge probability $1-\alpha$, a test set $\mathcal{D}_{\text{test}}$ and a method for predicting conformal sets $\hat{C}(\cdot)$, we resort to the following metrics.
Set $\text{Set size}(\mathcal{D}_{\text{test}})=\frac{1}{|\mathcal{D}_{\text{test}}|}\sum_{\mathbf{x}\in\mathcal{D}_{\text{test}}}|\hat{C}(\mathbf{x})|$ (7)
$\text{Cov}(\mathcal{D}_{\text{test}})=\frac{1}{|\mathcal{D}_{\text{test}}|}\sum_{\mathbf{x},y\in\mathcal{D}_{\text{test}}}\mathbbm{1}_{y\in\hat{C}(\mathbf{x})}$ (8)
$\text{CovGap}(\mathcal{D}_{\text{test}})=\frac{1}{|\mathcal{Y}|}\sum_{k\in\mathcal{Y}}\big{|}\text{Cov}(\mathcal{D}_{\text{test},k})-(1-\alpha)\big{|}$ (9)
$\text{MCCC}(D_{\text{test}})=\min_{k\in\mathcal{Y}}\big{(}\text{Cov}(\mathcal{D}_{\text{test},k})\big{)}$ (10)
Where Equation 7 is the average predicted set size across the test set (sometimes refered to as efficiency), Equation 8 is the empirical marginal coverage, Equation 9 represents the average gap between the empirical coverage and the class-conditioned coverage, and Equation 10 is the min class-conditioned coverage (MCCC).
B Models Used and Implementation Details
Foundation models.
We employ three different, and diverse foundation models for vision tasks, which differ in the number of modalities for training (i.e., vision vs vision-language), strategy used for training (i.e., contrastive vs self-supervised learning), and backbones used (i.e., vision transformers vs convolutional neural networks). More conctretely, the DINO [8, 40] models employed in this work are ViT-based models trained in a self-supervised manner. CLIP [46] is a vision-language model, which trains a vision and a text encoder in a contrastive fashion. Last, VICReg [3] are CNNs trained in a self-supervised manner using contrastive learning.
Implementation details.
All linear probing heads are trained with a cross-entropy loss with the Adam [25] optimizer and a learning rate of $10^{-4}$, whose patience is set to $10$. The implementation of the training is based on PyTorch [42], and the conformal predictions methods on TorchCP [63]. For RAPS, we used $k_{\text{RAPS}}=2$ and $\lambda_{\text{RAPS}}=0.1$ as hyperparameters. Across all experiments, we use $\alpha=0.1$, except for CIFAR-10 for which we use $\alpha=0.05$.
C Relationship Between Model Performance and CP Metrics
Table 4, Table 5, and Table 6 report the numerical values corresponding to Fig. 1, showcasing the relationship between the linear probing accuracy of each model, and the CP performance in terms of set size and coverage gap for ImageNet, CIFAR-10, and CIFAR-100, respectively.
D Model performance for the optimal temperature
Fig. 8, 9, 10, 11, 12, and 13 showcase the impact of temperature $T$ on the ECE, set size, minimum class-conditional coverage, and coverage gap across all foundation models. The observations on these figures strongly align with the findings on the main paper as, regardless of the model, APS yields optimal coverage gap close to the optimal temperature point. Indeed, looking at the different models (i.e., DINO-based, VICReg and CLIP), we can observe that the behaviour of RAPS (in terms of coverage gap), strongly varies across models. Additionally, Fig. 14 shows the evolution of $q_{\alpha}$, the threshold on the conformal scores $s$ before and after calibration. When applying temperature scaling, the distribution of softmax across classes approaches a uniform distribution, lowering the scores for the most likely classes. Intuitively, this means that more classes will need to be included in the set to ensure coverage, leading to a decrease in threhsold $q_{\alpha}$. This decrease is consistent with the observed increase in set size, which can be seen in Fig. 15, showing the difference in set size when applying temperature scaling for APS on CLIP. Note that this trend holds across models, as a similar behaviour for DINO is observed in Fig. 7 in the main paper. Note that both Fig. 7 and 15 only show cases where the set sizes are different between the calibrated and uncalibrated model.
Furthermore, we report in Table 7 a more complete version of the results presented in Table 2, containing all explored models.
E Conformal set size analysis
We show the difference between the conformal set size for APS when applied to $\text{ViT}_{\text{ImageNet}}$ and $\text{ViT}_{\text{DINO-S}}$ in Fig. 20 and for $\text{ViT}_{\text{ImageNet}}$ and $\text{ViT}_{\text{MetaCLIP}}$ in Fig. 19, similarly to what is shown in Fig. 4.
F Domain Shift
Table 8 reports the numerical values for Fig. 4, showcasing APS’s strong performance in terms of marginal and conditional coverage under distribution shift, at the cost of decreasing the set efficiency. Fig. 16, 17, and 18 depict the distribution of class-conditional coverage for DINOv2-B, VICReg (ResNet 50x2), and CLIP, respectively. The kernel density estimate plots show the class-conditional coverage for a distribution shift with ImageNet-A (top), ImageNet-R (middle), and ImageNet-V2 (bottom), for APS (left) and RAPS (right). These results clearly demonstrate the stronger resistance to distribution shift for APS compared to RAPS, even when both methods show a minimum class-conditional coverage of 0.
G Few-Shot Adaptation
We present in Table 9 and Table 11 a detailed version of Table 3 for 16 shots, where the OOD section (in Table 3) corresponds to the average across all ImageNet variants of Table 9, and the ID section corresponds to the average across all 11 datasets of Table 11. Table 10 shows in distribution results for adapting only 4 shots, where we observe slightly higher set size and coverage gap compare to the results for 16 shots.
AvgSize ($\downarrow$) MCCC ($\uparrow$)
Accuracy ($\uparrow$) LAC APS RAPS LAC APS RAPS
DINO-S 0.7492 3.27 10.02 4.23 0.072 0.059 0.074
DINO-B 0.7743 2.58 10.73 3.03 0.072 0.058 0.073
DINOv2-S 0.7948 1.87 6.66 2.12 0.073 0.057 0.071
DINOv2-B 0.8393 1.36 7.50 1.67 0.076 0.057 0.065
DINOv2-L 0.8624 1.21 6.77 1.50 0.079 0.057 0.063
DINOv2-G 0.8666 1.18 4.08 1.44 0.079 0.058 0.063
VICReg (ResNet-50) 0.7207 4.24 14.22 5.73 0.070 0.058 0.071
VICReg (ResNet-50x2) 0.7570 3.08 13.49 3.89 0.070 0.056 0.070
VICReg (ResNet-200x2) 0.7804 2.50 12.05 2.98 0.071 0.056 0.071
CLIP 0.7201 3.03 9.50 3.73 0.069 0.058 0.070
MetaCLIP 0.7580 2.39 9.43 2.84 0.479 0.535 0.467
Table 4: Linear probing accuracy, average set size, and minimum class-conditional coverage for all models on ImageNet.
AvgSize ($\downarrow$) MCCC ($\uparrow$)
Accuracy ($\uparrow$) LAC APS RAPS LAC APS RAPS
DINO-S 0.8711 1.36 1.68 1.53 0.023 0.010 0.017
DINO-B 0.9165 1.14 1.41 1.30 0.024 0.009 0.013
DINOv2-S 0.9140 1.13 1.41 1.29 0.024 0.009 0.011
DINOv2-B 0.9511 1.01 1.22 1.14 0.025 0.008 0.009
DINOv2-L 0.9848 0.95 1.03 1.01 0.020 0.008 0.009
DINOv2-G 0.9932 0.95 0.99 0.98 0.020 0.008 0.008
VICReg (ResNet-50) 0.8506 1.49 1.84 1.69 0.021 0.011 0.018
VICReg (ResNet-50x2) 0.8749 1.32 1.65 1.50 0.022 0.010 0.014
VICReg (ResNet-200x2) 0.8481 1.47 1.85 1.66 0.022 0.011 0.017
CLIP 0.8823 1.33 1.71 1.51 0.021 0.009 0.013
MetaCLIP 0.8926 1.22 1.59 1.41 0.903 0.933 0.928
Table 5: Linear probing accuracy, average set size, and minimum class-conditional coverage for all models on CIFAR-10.
AvgSize ($\downarrow$) MCCC ($\uparrow$)
Accuracy ($\uparrow$) LAC APS RAPS LAC APS RAPS
DINO-S 0.6668 3.46 5.64 4.29 0.050 0.037 0.055
DINO-B 0.7379 2.19 4.07 2.50 0.050 0.034 0.050
DINOv2-S 0.7515 2.06 4.06 2.40 0.049 0.035 0.050
DINOv2-B 0.8116 1.39 3.40 1.76 0.054 0.034 0.042
DINOv2-L 0.8885 1.02 2.37 1.35 0.061 0.035 0.036
DINOv2-G 0.9235 0.95 1.93 1.21 0.067 0.035 0.036
VICReg (ResNet-50) 0.6503 3.83 6.15 4.83 0.050 0.034 0.056
VICReg (ResNet-50x2) 0.6902 3.06 5.30 3.78 0.054 0.035 0.056
VICReg (ResNet-200x2) 0.6461 3.95 6.23 4.91 0.053 0.037 0.058
CLIP 0.6667 3.52 6.28 4.18 0.050 0.037 0.055
MetaCLIP 0.7015 2.79 5.25 3.27 0.683 0.769 0.652
Table 6: Linear probing accuracy, average set size, and minimum class-conditional coverage for all models on CIFAR-100.
(a) DINO-S
(b) DINO-B
(c) DINOv2-S
(d) DINOv2-B
(e) DINOv2-L
(f) DINOv2-G
Figure 8: Impact of the temperature $T$ on the ECE, set size, MCCC, and CovGap, for DINO-S, DINO-B, DINOv2-S, DINOv2-B, DINOv2-L, and DINOv2-G on ImageNet. Red vertical line indicates the uncalibrated model performance ($T=1$). Green vertical line indicates the model performance for the optimal temperature.
(a) VICReg (ResNet 50)
(b) VICReg (ResNet 50x2)
(c) VICReg (ResNet 200x2)
(d) CLIP
(e) MetaCLIP
Figure 9: Impact of the temperature $T$ on the ECE, set size, MCCC, and CovGap, for VICReg (ResNet 50), VICReg (ResNet 50x2), VICReg (ResNet 200x2), and CLIP on ImageNet. Red vertical line indicates the uncalibrated model performance ($T=1$). Green vertical line indicates the model performance for the optimal temperature.
(a) DINO-S
(b) DINO-B
(c) DINOv2-S
(d) DINOv2-B
(e) DINOv2-L
(f) DINOv2-G
Figure 10: Impact of the temperature $T$ on the ECE, set size, MCCC, and CovGap, for DINO-S, DINO-B, DINOv2-S, DINOv2-B, DINOv2-L, and DINOv2-G on CIFAR-10. Red vertical line indicates the uncalibrated model performance ($T=1$). Green vertical line indicates the model performance for the optimal temperature.
(a) VICReg (ResNet 50)
(b) VICReg (ResNet 50x2)
(c) VICReg (ResNet 200x2)
(d) CLIP
(e) MetaCLIP
Figure 11: Impact of the temperature $T$ on the ECE, set size, MCCC, and CovGap, for VICReg (ResNet 50), VICReg (ResNet 50x2), VICReg (ResNet 200x2), and CLIP on CIFAR-10. Red vertical line indicates the uncalibrated model performance ($T=1$). Green vertical line indicates the model performance for the optimal temperature.
(a) DINO-S
(b) DINO-B
(c) DINOv2-S
(d) DINOv2-B
(e) DINOv2-L
(f) DINOv2-G
Figure 12: Impact of the temperature $T$ on the ECE, set size, MCCC, and CovGap, for DINO-S, DINO-B, DINOv2-S, DINOv2-B, DINOv2-L, and DINOv2-G on CIFAR-100. Red vertical line indicates the uncalibrated model performance ($T=1$). Green vertical line indicates the model performance for the optimal temperature.
(a) VICReg (ResNet 50)
(b) VICReg (ResNet 50x2)
(c) VICReg (ResNet 200x2)
(d) CLIP
(e) MetaCLIP
Figure 13: Impact of the temperature $T$ on the ECE, set size, MCCC, and CovGap, for VICReg (ResNet 50), VICReg (ResNet 50x2), VICReg (ResNet 200x2), and CLIP on CIFAR-100. Red vertical line indicates the uncalibrated model performance ($T=1$). Green vertical line indicates the model performance for the optimal temperature.
(a) LAC
(b) APS
(c) RAPS
Figure 14: ECE and average $q_{\alpha}$ threshold before and after calibration for (a) LAC, (b) APS, and (c) RAPS on ImageNet. Each pair represents one model before (square) and after (circle) calibration.
Figure 15: Difference in conformal set size (i.e., efficiency) when applying temperature scaling on APS and CLIP ($T=1.1$).
ECE ($\downarrow$) AvgSize ($\downarrow$) MCCC ($\uparrow$)
(×$\times 10^{-2}$) LAC APS RAPS LAC APS RAPS
DINO-S 5.59/3.02 3.27/3.26 10.02/11.78 4.23/3.99 0.433/0.427 0.477/0.480 0.412/0.415
DINO-B 3.94/2.12 2.58/2.55 10.73/13.14 3.03/3.03 0.416/0.398 0.521/0.517 0.391/0.409
DINOv2-S 2.71/1.37 1.87/1.87 6.66/8.34 2.12/2.19 0.370/0.351 0.520/0.538 0.384/0.409
DINOv2-B 2.85/1.81 1.36/1.36 7.50/10.46 1.67/1.77 0.414/0.406 0.476/0.482 0.489/0.498
DINOv2-L 2.99/1.81 1.21/1.21 6.77/9.67 1.50/1.59 0.326/0.336 0.502/0.522 0.453/0.487
DINOv2-G 3.66/2.10 1.18/1.18 4.08/5.69 1.44/1.50 0.244/0.243 0.458/0.484 0.350/0.375
VICReg (RN-50) 3.12/1.90 4.24/4.19 14.22/16.6 5.73/5.49 0.472/0.468 0.566/0.570 0.453/0.456
VICReg (RN-50x2) 2.34/2.16 3.08/3.09 13.49/16.06 3.89/3.80 0.442/0.430 0.549/0.557 0.445/0.446
VICReg (RN-200x2) 2.21/1.90 2.50/2.49 12.05/14.61 2.98/3.01 0.440/0.430 0.567/0.566 0.462/0.471
CLIP 2.70/1.63 3.03/3.01 9.50/11.11 3.73/3.53 0.434/0.441 0.556/0.564 0.418/0.414
MetaCLIP 2.63/1.99 2.39/2.40 9.43/11.31 2.84/2.84 0.479/0.477 0.535/0.541 0.467/0.469
Table 7: Quantitative impact of calibrating vision foundation models. Average size and minimum class conditional coverage are reported between uncalibrated/calibrated models ($T=1$ and $T=1.1$ respectively) for several models, with ImageNet as benchmark dataset.
Set size ($\downarrow$) Coverage CovGap ($\downarrow$)
LAC APS RAPS LAC APS RAPS LAC APS RAPS
ImageNet DINOv2-B 1.36 7.50 1.67 0.900 0.900 0.900 0.414 0.476 0.489
VICReg (RN 50x2) 3.08 13.49 3.89 0.900 0.900 0.900 0.442 0.549 0.445
CLIP 3.03 9.50 3.73 0.900 0.901 0.900 0.434 0.556 0.418
ImageNet-V2 DINOv2-B 1.38 9.47 1.74 0.879 0.892 0.881 0.081 0.187 0.068
VICReg (RN 50x2) 3.15 12.54 3.89 0.867 0.871 0.870 0.130 0.175 0.143
CLIP 2.88 6.79 3.58 0.877 0.864 0.882 0.023 0.034 0.029
ImageNet-Sketch DINOv2-B 1.35 10.90 1.91 0.884 0.904 0.893 0.221 0.322 0.275
VICReg (RN 50x2) 3.49 21.98 4.51 0.826 0.879 0.830 0.119 0.223 0.114
CLIP 3.16 12.98 3.99 0.877 0.900 0.878 0.182 0.293 0.185
ImageNet-A DINOv2-B 1.63 25.30 3.55 0.646 0.897 0.734 0.000 0.000 0.000
VICReg (RN 50x2) 8.43 59.83 7.82 0.531 0.879 0.511 0.000 0.000 0.000
CLIP 6.79 39.42 6.45 0.648 0.912 0.624 0.000 0.000 0.000
ImageNet-R DINOv2-B 1.31 7.49 1.94 0.867 0.910 0.879 0.282 0.532 0.408
VICReg (RN 50x2) 5.84 27.36 5.65 0.761 0.897 0.742 0.198 0.429 0.186
CLIP 3.38 11.93 4.13 0.893 0.930 0.890 0.422 0.604 0.386
Table 8: Evaluation under domain-shift. Set size($\downarrow$), coverage($\uparrow$), and MCCC($\uparrow$) across three CP methods and three foundation models.
(a) DINOv2-B, ImageNet-A
(b) DINOv2-B, ImageNet-R
(c) DINOv2-B, ImageNet-Sketch
Figure 16: Domain shift analysis. Distribution of class-conditional coverages for DINOv2-B on ImageNet-A, ImageNet-R, and ImageNet-Sketch: APS (left) and RAPS (right).
(a) VICReg (RN 50x2), ImageNet-A
(b) VICReg (RN 50x2), ImageNet-R
(c) VICReg (RN 50x2), ImageNet-Sketch
Figure 17: Domain shift analysis. Distribution of class-conditional coverages for VICReg (RN 50x2) on ImageNet-A, ImageNet-R, and ImageNet-Sketch: APS (left) and RAPS (right).
(a) CLIP, ImageNet-A
(b) CLIP, ImageNet-R
(c) CLIP, ImageNet-Sketch
Figure 18: Domain shift analysis. Distribution of class-conditional coverages for CLIP on ImageNet-A, ImageNet-R, and ImageNet-Sketch: APS (left) and RAPS (right).
Figure 19: $\text{ViT}_{\text{ImageNet}}$ vs $\text{ViT}_{\text{MetaCLIP}}$. Analyzing the difference in set size for APS between a $\text{ViT}_{\text{MetaCLIP}}$ and $\text{ViT}_{\text{ImageNet}}$. Equal set sizes not shown.
Figure 20: $\text{ViT}_{\text{ImageNet}}$ vs $\text{ViT}_{\text{DINO-S}}$. Analyzing the difference in set size for APS between a $\text{ViT}_{\text{DINO-S}}$ and $\text{ViT}_{\text{ImageNet}}$. Equal set sizes not shown.
Set size ($\downarrow$) CovGap ($\downarrow$)
LAC APS RAPS LAC APS RAPS
ImageNet ZS 2.81 10.07 3.23 0.086 0.069 0.084
ZSLP 2.17 6.60 2.45 0.076 0.060 0.074
CLAP [53] 2.15 6.83 2.45 0.077 0.060 0.074
CoOp [70] 2.44 6.32 2.79 0.080 0.066 0.080
KgCoOp [17] 2.47 7.18 2.84 0.081 0.065 0.081
-V2 ZS 4.64 17.34 5.72 0.128 0.124 0.128
ZSLP 3.89 12.51 4.81 0.126 0.121 0.125
CLAP [53] 3.72 12.72 4.62 0.126 0.122 0.125
CoOp [70] 4.08 11.26 4.87 0.126 0.123 0.126
KgCoOp [17] 4.09 12.65 4.93 0.126 0.123 0.126
-Sketch ZS 14.74 37.18 20.39 0.099 0.090 0.099
ZSLP 16.46 37.27 22.79 0.098 0.091 0.098
CLAP [53] 15.20 34.90 21.29 0.098 0.090 0.098
CoOp [70] 16.06 34.36 22.12 0.098 0.091 0.099
KgCoOp [17] 15.46 34.35 21.74 0.098 0.092 0.099
-A ZS 9.42 16.59 11.28 0.094 0.085 0.095
ZSLP 10.92 18.63 12.47 0.087 0.080 0.088
CLAP [53] 9.36 16.56 11.38 0.087 0.082 0.087
CoOp [70] 9.23 14.65 10.72 0.087 0.084 0.090
KgCoOp [17] 8.66 14.39 10.19 0.093 0.084 0.096
-R ZS 1.92 5.77 2.31 0.059 0.039 0.054
ZSLP 2.07 5.62 2.36 0.058 0.041 0.056
CLAP [53] 1.92 5.46 2.26 0.059 0.039 0.053
CoOp [70] 2.06 5.11 2.40 0.061 0.042 0.063
KgCoOp [17] 1.96 5.32 2.29 0.060 0.041 0.058
Table 9: Set size and CovGap for CLIP (ViT-B backbone) on ImageNet and its variants.
Set size ($\downarrow$) CovGap ($\downarrow$)
LAC APS RAPS LAC APS RAPS
Aircraft ZS 18.7 19.75 20.74 0.138 0.139 0.127
ZSLP 9.73 11.13 11.55 0.082 0.073 0.086
CoOp [70] 9.41 10.62 11.38 0.081 0.076 0.088
KgCoOp [17] 9.46 10.82 11.0 0.084 0.076 0.086
CLAP [53] 9.73 11.0 11.51 0.082 0.073 0.086
Caltech-101 ZS 0.94 1.95 1.20 0.136 0.092 0.092
ZSLP 0.91 1.20 1.05 0.147 0.085 0.085
CoOp [70] 0.92 1.14 1.06 0.143 0.085 0.087
KgCoOp [17] 0.92 1.25 1.08 0.147 0.084 0.085
CLAP [53] 0.91 1.31 1.08 0.144 0.085 0.086
DTD ZS 11.78 13.3 13.66 0.122 0.117 0.121
ZSLP 4.74 6.83 5.67 0.092 0.077 0.091
CoOp [70] 5.68 6.46 7.00 0.099 0.093 0.098
KgCoOp [17] 5.33 6.22 6.33 0.097 0.083 0.096
CLAP [53] 4.75 6.83 5.68 0.094 0.079 0.093
EuroSAT ZS 4.55 4.78 5.06 0.103 0.101 0.102
ZSLP 1.62 2.22 1.99 0.043 0.051 0.042
CoOp [70] 1.82 2.16 1.95 0.043 0.028 0.034
KgCoOp [17] 1.51 1.92 1.76 0.028 0.030 0.032
CLAP [53] 1.73 2.33 2.06 0.043 0.048 0.041
Flowers-102 ZS 9.26 9.52 15.15 0.180 0.175 0.179
ZSLP 1.09 2.36 1.65 0.132 0.089 0.095
CoOp [70] 1.13 1.80 1.51 0.104 0.080 0.083
KgCoOp [17] 1.19 2.04 1.64 0.114 0.082 0.085
CLAP [53] 1.15 2.47 1.70 0.137 0.092 0.097
Food-101 ZS 1.14 1.89 1.45 0.054 0.026 0.031
ZSLP 1.11 1.82 1.42 0.046 0.023 0.026
CoOp [70] 1.15 1.77 1.41 0.050 0.026 0.030
KgCoOp [17] 1.10 1.86 1.43 0.050 0.025 0.027
CLAP [53] 1.10 1.83 1.42 0.047 0.023 0.026
ImageNet ZS 2.81 10.07 3.23 0.086 0.069 0.084
ZSLP 2.48 8.02 2.88 0.078 0.062 0.078
CoOp [70] 2.69 7.33 3.03 0.081 0.067 0.081
KgCoOp [17] 2.65 8.60 3.01 0.082 0.066 0.082
CLAP [53] 2.44 8.05 2.82 0.079 0.062 0.079
Oxford Pets ZS 1.05 1.43 1.33 0.109 0.067 0.072
ZSLP 0.95 1.26 1.19 0.080 0.039 0.038
CoOp [70] 0.95 1.17 1.13 0.084 0.036 0.036
KgCoOp [17] 0.94 1.20 1.15 0.079 0.038 0.037
CLAP [53] 0.95 1.28 1.20 0.084 0.039 0.038
Stanford Cars ZS 2.24 3.09 2.49 0.109 0.080 0.102
ZSLP 1.62 2.39 1.94 0.083 0.061 0.067
CoOp [70] 1.95 2.64 2.21 0.087 0.065 0.078
KgCoOp [17] 1.99 2.77 2.29 0.096 0.069 0.084
CLAP [53] 1.63 2.40 1.96 0.085 0.061 0.067
SUN397 ZS 2.83 5.80 3.21 0.093 0.073 0.093
ZSLP 2.00 3.96 2.29 0.080 0.060 0.072
CoOp [70] 2.35 3.76 2.64 0.082 0.068 0.081
KgCoOp [17] 2.34 4.18 2.60 0.083 0.065 0.082
CLAP [53] 2.01 3.99 2.31 0.080 0.060 0.072
UCF101 ZS 2.89 5.24 3.83 0.124 0.096 0.124
ZSLP 1.58 3.12 1.98 0.112 0.074 0.088
CoOp [70] 1.97 3.12 2.18 0.114 0.077 0.110
KgCoOp [17] 1.74 3.06 2.06 0.114 0.070 0.096
CLAP [53] 1.63 3.25 2.00 0.114 0.076 0.090
Table 10: Set size and CovGap for CLIP (ViT-B backbone) on 11 datasets across several few-shot adaptation approaches. Training for 4 shots.
Set size ($\downarrow$) CovGap ($\downarrow$)
LAC APS RAPS LAC APS RAPS
Aircraft ZS 18.7 19.75 20.74 0.138 0.139 0.127
ZSLP 8.14 9.56 9.43 0.079 0.069 0.082
CLAP [53] 8.20 9.59 9.43 0.079 0.070 0.083
CoOp [70] 7.19 8.09 8.88 0.080 0.070 0.084
KgCoOp [17] 7.53 8.72 9.13 0.080 0.075 0.084
Caltech-101 ZS 0.94 1.95 1.20 0.136 0.092 0.092
ZSLP 0.91 1.12 1.02 0.151 0.083 0.082
CLAP [53] 0.91 1.22 1.06 0.146 0.083 0.085
CoOp [70] 0.91 1.06 1.02 0.148 0.084 0.085
KgCoOp [17] 0.91 1.19 1.06 0.150 0.082 0.083
DTD ZS 11.78 13.3 13.66 0.122 0.117 0.121
ZSLP 2.91 4.90 3.34 0.081 0.071 0.073
CLAP [53] 3.03 5.12 3.48 0.080 0.069 0.079
CoOp [70] 3.37 4.50 3.94 0.084 0.076 0.081
KgCoOp [17] 3.53 4.81 4.04 0.082 0.073 0.082
EuroSAT ZS 4.55 4.78 5.06 0.103 0.101 0.102
ZSLP 1.46 2.04 1.84 0.053 0.045 0.038
CLAP [53] 1.53 2.09 1.90 0.051 0.044 0.038
CoOp [70] 1.23 1.59 1.50 0.041 0.026 0.031
KgCoOp [17] 1.28 1.73 1.61 0.045 0.032 0.032
Flowers-102 ZS 9.26 9.52 15.15 0.180 0.175 0.179
ZSLP 1.03 2.19 1.58 0.132 0.087 0.091
CLAP [53] 1.09 2.30 1.65 0.135 0.089 0.094
CoOp [70] 0.93 1.30 1.17 0.094 0.078 0.078
KgCoOp [17] 0.93 1.44 1.24 0.102 0.079 0.078
Food-101 ZS 1.14 1.89 1.45 0.054 0.026 0.031
ZSLP 1.08 1.77 1.39 0.044 0.023 0.025
CLAP [53] 1.07 1.79 1.39 0.045 0.023 0.024
CoOp [70] 1.11 1.64 1.35 0.050 0.025 0.03
KgCoOp [17] 1.08 1.75 1.38 0.048 0.022 0.026
ImageNet ZS 2.81 10.07 3.23 0.086 0.069 0.084
ZSLP 2.17 6.60 2.44 0.076 0.060 0.074
CLAP [53] 2.15 6.85 2.45 0.077 0.060 0.074
CoOp [70] 2.39 5.27 2.73 0.078 0.064 0.077
KgCoOp [17] 2.30 5.95 2.64 0.078 0.062 0.077
Oxford Pets ZS 1.05 1.43 1.33 0.109 0.067 0.072
ZSLP 0.94 1.24 1.17 0.081 0.040 0.037
CLAP [53] 0.95 1.27 1.19 0.089 0.039 0.038
CoOp [70] 0.94 1.16 1.13 0.083 0.040 0.041
KgCoOp [17] 0.94 1.20 1.15 0.083 0.038 0.038
Stanford Cars ZS 2.24 3.09 2.49 0.109 0.08 0.102
ZSLP 1.27 1.98 1.66 0.082 0.058 0.057
CLAP [53] 1.32 2.04 1.7 0.080 0.057 0.057
CoOp [70] 1.37 1.94 1.66 0.078 0.057 0.059
KgCoOp [17] 1.39 2.04 1.73 0.080 0.059 0.061
SUN397 ZS 2.83 5.80 3.21 0.093 0.073 0.093
ZSLP 1.72 3.49 2.09 0.075 0.055 0.063
CLAP [53] 1.76 3.55 2.12 0.075 0.056 0.064
CoOp [70] 1.87 2.73 2.04 0.073 0.060 0.072
KgCoOp [17] 1.85 2.91 2.07 0.076 0.059 0.070
UCF101 ZS 2.89 5.24 3.83 0.124 0.096 0.124
ZSLP 1.36 2.85 1.82 0.106 0.068 0.077
CLAP [53] 1.41 2.94 1.87 0.108 0.067 0.079
CoOp [70] 1.51 2.28 1.76 0.108 0.068 0.095
KgCoOp [17] 1.41 2.47 1.77 0.106 0.067 0.081
Table 11: Set size and CovGap for CLIP (ViT-B backbone) on 11 datasets across several few-shot adaptation approaches. Training for 16 shots.